### v1

# Adaptive behavior with stable synapses

Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters. However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could also be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks. We argue that such a framework can describe the psychometry of context-dependent tasks on humans and other species, solving the incoherence of plasticity timescales. When the context is changed, the network is dynamically reconfigured, and the predicted output undergoes dynamic updates until it aligns with the information embedded in the context. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
Introduction
The study of behavioral adaptation, observed in both humans and animals, has been a longstanding subject of research. The rapidity with which behaviors can change in response to new cues challenges the conventional explanation of synaptic plasticity. This led researchers to explore mechanisms that underlie such flexible adaptations [1].
In-context learning as an emerging property in AI
In-context learning (ICL) is the capacity of a model to adapt its behavior, without weight updates, to solve tasks not encountered during training. Initially, ICL was observed in architectures tailored for few-shot learning [2] or even zero-shot learning [3]. However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [4, 5]. Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[6, 7, 8]. Despite some intriguing explanations [9, 10, 11], the emergence of this property remains not completely understood. In particular, it is not clear how to associate it with in-context learning abilities observed in biological networks. The major contribution of this work is the definition of a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
Limitations of transformers architecture
Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length. This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding. To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs. Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks. Zucchet et al. [14] explore the efficacy of these deep linear RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and facilitating in-context learning.
Biological support for in-context learning
Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers.
Recent findings on dendritic computational properties [15] and on the complexity of pyramidal neurons dynamics [16] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [17, 18, 19, 20]. It has has been proposed that segregation of dendritic input [18] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [20] would support backpropagation in biological neurons. In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network.
Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of stationary and temporal datasets, as well as for reinforcement learning.
This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
Methods
Dynamical adaptation
Consider a generic learning task, in which the response $y$ to a state $x$ and is influenced by a set of parameters $w$. The latter are adjusted in function of an internal error $e$. This can be rewritten in a generic formulation as follows:
$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f)\\ \tau_{y}\,\dot{y}&=Y(y,w,x)\\ \tau_{w}\,\dot{w}&=W(w,x,e)\end{cases}$ (1)
where $f$ is feedback from the environment (e.g. the reward, the target behavior, …). Alternatively to the standard interpretation of $w$ as synaptic weights of a neuronal network, we consider them as the state variables of a dynamical system coupled to the dynamics of $y$, expressed by the activity of an auxiliary network. For this reason, we also refer to them as “virtual” weights.
Dynamical adaptation for linear regression
Consider now the task of learning a linear relationship between input $x$ and output $y$. In the conventional approach, a neural network is initialized with random parameters, essentially making random guesses. The model estimation can be improved through iterative adjustments of the network parameters (e.g. the synaptic weights), following the gradient of errors evaluated on a training dataset $(x_{i}^{data},y_{i}^{data})$. One step of the gradient would update the weights of a linear model $y=wx$, following $\delta w=-\sum_{i}(y-y_{i}^{data})x^{data}_{i}$, where $\eta$ is the learning rate. By the end of the training process, the network aims at accurately predict the output $y^{test}$ for a new input $x^{test}$. An equivalent error reduction can be achieved by updating the target output $y_{i}^{data}$ to $y_{i}^{data}-wx_{i}^{data}=y_{i}^{data}-y_{i}^{train}$ (let’s denote this the latter quantity as $e_{i}$), where $y_{i}^{train}=wx_{i}^{data}$ is the current model estimation of $y_{i}^{data}$. $e_{i}$ can then be updated iteratively, and when it reaches the value zero, it indicates that the interpretation of the dataset is coherent, and the network predictions on new data $x^{test}$ are reliable. We can interpret this dynamics as changing the way $x^{test}$ is processed in function of the virtual weights $w$. Typically, when a network dynamically alters the way it receives an input, it is referred to as an attention mechanism. Following this concept, our goal is to project the dataset $(x_{i}^{data},y_{i}^{data})$ onto the network, influencing the way a new input $x^{test}$ (projected simultaneously) is processed. The information that drives and influences the attention mechanisms, $(x_{i}^{data},y_{i}^{data})$ in this case, is usually referred to as a context. When the input vector is one dimensional, the task can be solved by a network composed of three modules, one evaluating the error $e$, one evaluating the necessary update $\delta w$, and one evaluating the response $y$ to input data (see Fig. 1A). Their dynamics can be explicitly written as follows  [^1] Here we model the $y$ to be an instantaneous function of the variables $w,x$. This is equivalent to considering $Y(y,w,x)=xw-y$ in Equation (1) and taking the limit $\tau_{y}\to 0$. Similar considerations hold for $e$.:
$\begin{cases}e_{i}&=(y^{data}_{i}-x_{i}^{data}w)\\ \tau_{w}\,\dot{w}&=\sum_{i=1}^{M}x_{i}^{data}\,e_{i}\\ y&=x^{test}w\end{cases}$ (2)
with $e_{i}(0)=y_{i}^{data}-w_{0}x_{i}^{data}$, and $w_{0}$ are the initial weights. The conditioning dataset is $(x_{i}^{data},y_{i}^{data})$, with $i=1,...M$, while $(x^{test},y^{test})$ is one element of the test set. In this case one example ($M=1$) is sufficient to determine the response to new inputs. This formulation is equivalent to what derived in [11], with the difference that here many iteration are recursively evaluated over time, allowing the convergence to the gradient descent solution. See Results section for examples of applications. This approach can be extended to multidimensional linear datasets (see Supporting Information).
Dynamical supervised learning for temporal trajectory
The above formulation is well-posed for stationary dataset, however such a framework (the temporal convergence of the error to zero and of the response to the desired value) is tricky for temporal trajectories, which requires a little change of paradigm, to achieve a formulation that holds at each time step.
First of all we introduce a target temporal trajectory $y^{targ}(t)$, and define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset. In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$. We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$.
This together with Eq. (2) allows us to write:
$\begin{cases}e&=(y^{targ}-y)\\ y&=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{x}\,e\end{cases}$ (3)
where we removed the dependence on time $t$ for simplicity. The operations required are nonlinear, and usually are naturally implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights. However, here $\bm{w}$ are not actual weights but rather dynamical variables.
We propose, as possible implementation of this, that such non-linear functions are computed by two neural networks networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{x},e)\simeq W(\bm{x},e)=\bm{x}\,e\end{cases}$ (4)
In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm).
In addition, we introduce a novel concept known as a gain-modulated network. This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]). These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions. More details in the section "Gain Modulated Reservoir Computing". We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents.
The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$. However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments.
We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions. For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$. This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above. The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time. This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
$\begin{cases}\bm{e}&=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{e}\odot\bm{x}\end{cases}$ (5)
This is the dynamics obtained by evaluating the policy gradient with respect to the virtual weights $\bm{w}$. $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones. $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [23, 24]) at time $t$. It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
This formulation holds for a null discount factor [23, 24]). See Supplementary information for a complete derivation and for the description of the general case. The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (6)
Algorithm Distillation
To demonstrate the ability of this architecture to learn in context, we train networks Y and W using an Algorithmic Distillation protocol, as defined in [25]. Namely, we consider a family of environments $\mathcal{A}$ with different reward distributions  [^2] A family of environments is a set of possible environments with the same action-state transitions but different reward distributions.. We then train the aforementioned networks using $N_{s}$ learning histories obtained through policy gradient methods from a subset of these environments, designated as $\mathcal{A}_{\text{ID}}$. The network performance is then tested out of distribution, i.e., on the rest of the environments $\mathcal{A}_{\text{OOD}}:=\mathcal{A}\setminus\mathcal{A}_{\text{ID}}$
Gain modulated reservoir computing (GM-RC)
Reservoir Computing
Reservoir Computing (RC) represents a paradigm in machine learning that provides a model for biologically plausible computation and learning, drawing inspiration from the information processing mechanisms observed in biological neural networks.
In this approach, a random RNN is employed to extract features from a time-dependent signal $\bm{x}(t)$. The dynamics of the RNN describe the evolution of $N$ hidden units $\bm{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
$\tau_{z}\dot{\bm{z}}=\phi\left(J\bm{z}+R\bm{x}\right)-\bm{z}$ (7)
The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([26]). Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively. Subsequently, the features extracted by the network can be utilized to predict a target signal $\bm{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\bm{y}^{targ}(t)-\Theta\bm{z}(t)|^{2}_{2}$ is minimized. The reservoir is then implementing a map
$\bm{y}(t)=\mathrm{RNN}_{\Theta}(\{\bm{x}(s)\}_{s\leq t})$ (8)
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies. 333 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit. In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\bm{x}\right)$. Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights. This architecture is also referred to in the literature as the Extreme Learning Machine [28].
Gain modulated network architecture
Analysing Equations (4) and (6) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters. Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC). This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [22, 29]. Here, we consider an additional input source $\bm{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$. Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope. Consequently, the resulting RNN equation is formulated as:
$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+R\bm{x})\right)-\bm{z}$ (9)
Where $\bm{b}^{\text{ap}}$ is a constant bias vector. Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the effect of the apical inputs on the network. Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation. The choice for the name $\bm{x}^{\text{ap}}$ is inspired by to current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high level information [22].
Results
Figure 1: Dynamical adaptation to changing context: A. Network architecture composed of 3 modules: $e$, $\delta w$ and $y$. The estimation on new data is produced by $y$, and is modulated by the error variable $e$ through $\delta w$. B. Example of the time course for the three variables in a single realization of the task. The values of $y$ converge to the correct values of (black dashed lines). C. This model accounts for the recovery of optimal performances (bottom, observed in humans and monkeys data) when the context suddenly changes, compared to a slower increase of performances achieved through synaptic plasticity.
Dynamical adaptation replaces synaptic plasticity
To demonstrate that suited responses can be provided relying on dynamical adaptation, we first consider a task requesting to learn a linear relationship between input $x^{data}$ and output $y^{data}$. As described in the Methods we consider a three module model (see Fig. 1A) described by the variables, $e$, $\delta w$, and $y$, whose dynamics is defined by Eq. (2), that is equivalent to following a gradient dynamics as respect to the parameters $w_{0}$. This update can be realized by treating $e$ as an input that, conditioned by the dataset $(x^{data},y^{data})$, evolves over time.
When $e\to 0$ the interpretation of the dataset is coherent, and the input-output mapping performed by the network on the new data is reliable. In this case one example is sufficient to determine the response to new inputs.
Fig. 1B shows the dynamics of three variables: the absolute value of $e$, and $\delta w$ as a consequence the error, are reduced over time, while the $y$ converges to the correct $y^{test}$ given the input $x^{test}$ (shown for 10 different new values).
We show the performances of the network whose parameters $w_{0}$ are trained through a gradient based plasticity on the rule $y^{data}=x^{data}$ (see first bar in the plots Fig. 1C), after the rule is switched to $y^{data}=-x^{data}$ (see, the second bar in which performances drop in both panels). In the top panel $w_{0}$ parameters are trained through gradient updates (we refer to Github repository https://github.com/cristianocapone/ABSS for details on parameters and implementation). In the bottom panel weights are not re-trained, but the last example (that is coherent with the new rule) is projected as a context (e.g. $(x^{data},y^{data})=(-1,1)$, starting the trial numbered as 1). This represents the agent understanding that the rule has been switched, and allows rapid and dynamic network reconfiguration and recovery of performances to optimal values, similarly to what observed in monkeys and humans data. An interesting example is the Wisconsin Card Sorting Test, in which participants are required to sort a deck of cards based on changing rules [30]. Initially, participants are given a rule (e.g., choosing by color) , and after a certain number of trials, the rule is changed without warning (e.g., choosing by shape). Participants then need to adapt to the new rule and adjust their strategy accordingly. The quick adaptation observed in this setup [31] reflects the ability to update mental models of the task, and flexibly apply new rules or strategies (similarly to Fig. 1C-bottom). On the other hand in some pathological subject, the adaptation after rule change appears to be slower [31], suggesting a learning process compatible with gradient-based plasticity (Fig. 1C-top).
Figure 2: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of 10 units (described by the vector $\bm{z}(t)$, depicted in pink) is utilized for learning a periodic trajectory, following the prescription of reservoir computing,. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios.
Dynamical adaptation for temporal trajectories
We aim to demonstrate that such an approach can be generalized to non-stationary scenarios.
We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm, we employ an RNN (Fig. 2A, pink), whose dynamics follows Equation (7), to extract temporal features $\bm{z}(t)$. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 2B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN. The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},\bm{w})\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\,e\end{cases}$ (10)
Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.2A green and orange respectively). Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section. These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$. We refer to Supporting Information for further details on the training procedure.
In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN. This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 2B).
In Fig.2B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line.
The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.2C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C).
Figure 3: Dynamical reinforcement learning: Multi Armed Bandits. A. Schematic of network architecture and task. Virtual weighs parameterize the agent’s policy (orange). At each round, an action (red) is sampled from the policy and played. A reward is then sampled from the current environment. A GM-network (green) then predicts the virtual parameter update. B. Comparing in distribution (ID) and out-of-distribution (OOD) expected reward at the 1000-th round for the policy gradient (PG) training source, reward gain-modulated network ($\gamma>0$), and a network without gain modulation $\gamma=0$. C. ID and OOD model performance varying gain modulation strength ($\gamma$). Solid lines indicate the median score (expected reward), computed over 20 independently trained models. The filled area indicates 25-75% confidence interval. D. Same as B but comparing OOD performance for different network sizes (number of hidden units). E. Robustness to hyperparameter variations in architectures with $(\gamma=0.9)$ and without ($\gamma=0.0$) gain modulation. We report the percentage of models that fail to generalize to OOD environments, varying the standard deviation of apical ($\sigma_{a}$) and basal ($\sigma_{b}$) random input weights. F. Regret comparison between the distilled and the original policy gradient algorithm. We report average regret per round in ID and OOD setting, compared with the one achieved by the PG training source.
Multi Armed Bandits
Dynamical adaptation is now investigated within the framework of reinforcement learning. We provide robust evidence supporting the hypothesis that gain modulation represents a fundamental component in implementing in-context learning in biological agents. We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [32]. Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability. During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 3A, green). This serves as an ideal test bed for evaluating the network ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action, as described in the Methods section. We systematically investigate the impact of the gain modulation on out-of-distribution performance, varying the parameter $\gamma$. Notably, when $\gamma=0$ the reward does not modulate the network gain.
We present our experimental findings in Fig. 3. Initially, we observe that integrating reward gain modulation into the network enables the agent to consistently converge to the optimal policy, even in out-of-distribution environments (see Fig.3B). To further explore this phenomenon, we trained multiple models while varying the strength of the modulation parameter ($\gamma$). We observe that performance improves with higher values of $\gamma$, allowing the generalization of learned rules out-of-distribution (Fig. 3C) and showcasing increased robustness to variations in hyperparameters (Fig. 3E). Furthermore, we observe that in the absence of gain modulation, the network is unable to generalize, merely memorizing the rules from the training set. This is evidenced by its persistence in selecting even-numbered arms even when tested in out-of-distribution environments where high rewards are allocated to odd-numbered arms (see Fig. 3E). Comparing the performances of models of different sizes (number $N$ of hidden units), we see that models with more reward gain modulation need significantly fewer neurons to achieve maximum scores in OOD environments (see Fig.3D).
Analysing the regret curves displayed in Fig. 3F we observe that a model with gain modulation learns a more data-efficient algorithm than its source even in OOD environments. On the contrary, a model without gain modulation fails to generalize to OOD environments, not converging to zero regret (see Fig. 3F, right panel).
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach unobservable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). E. Reward plotted against the number of trials, for the policy gradient (PG) algorithm (displayed in blue lines) applied in the dynamical PG scenario on both training positions (depicted in the left panel) and new positions (shown in the right panel). Thin line: single position, thick line: average across different positions. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [33]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility.
The horizontal and vertical positions of the agent are encoded separately using five input units $\bm{x}$ each, employing Gaussian activation functions with equally spaced centers ($0.4$ apart) and a width of $0.4$.
To achieve this task we consider a network architecture composed of two networks. One GM-network, , $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 4A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig. 4A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equations:
$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq\bm{x}\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\end{cases}$ (11)
Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig. 4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig. 4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions). After multiple trials, the agent successfully achieved precise targeting of the circles. Data collected from these experiments were utilized to train our networks to estimate gradients and scalar products, as defined in Eq. (11).
To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig. 4C, left panel) and new test positions (OOD, Fig. 4C, right panel). For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training. The agent’s precision closely matches that of the plastic policy gradient learning rule. We present the reward plotted against the number of trials for both training (Fig. 4E, right panel, pink lines) and test (Fig. 4C, right panel, pink lines) food locations.
We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig. 4B, D, F). Notably, while performances for ID food locations are acceptable (Fig. 4E, left panel, pink lines), those for OOD cases are extremely low (Fig. 4E, right panel, pink lines). This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
Discussion
It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 34], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8]. Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers. One intriguing explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11]. However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals. In this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
When considering a stationary dataset our model dynamically adapt to the context provided by the data, allowing predictions on new data, without synaptic plasticity. Remarkably, when the learning rule shifts abruptly, the network swiftly adapts without retraining, mimicking human cognitive flexibility observed in tasks like the Wisconsin Card Sorting Test. Conversely, when using gradient-based plasticity, the learning speed resembles what observed in certain pathological conditions [31]. When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward. We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates. Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response. We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
In conclusion, our approach provide an explicit framework to induce ICL in biologically plausible network, possibly opening the route to a formal understanding of ICL in biological agents.
Acknowledgments
This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’) to MM. LF is supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.
Source code availability
The source code is available under CC-BY license in the https://github.com/cristianocapone/ABSS public repository.

### v2

@@
 Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters.
-Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could also be supported by input segregation and dendritic amplification, features extensively observed in biological networks.
+Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks.
 We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks.
-We argue that such a framework can describe the psychometry of context-dependent tasks on humans and other species, solving the incoherence of plasticity timescales.
-When the context is changed, the network is dynamically reconfigured, and the predicted output undergoes dynamic updates until it aligns with the information embedded in the context.
 We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning.
 Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
-Introduction
+1 Introduction
 The study of behavioral adaptation, observed in both humans and animals, has been a longstanding subject of research.
@@
 Zucchet et al.
-[14] explore the efficacy of these deep linear RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and facilitating in-context learning.
+[14] explore the efficacy of these deep linear gated RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and implementing in-context supervised learning.
 Biological support for in-context learning
@@
 Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights.
-We present results for supervised learning of stationary and temporal datasets, as well as for reinforcement learning.
-This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
-Methods
-Dynamical adaptation
+We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations.
+This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behaviour in both natural and artificial intelligence domains.
+Our work generalizes and provides a biologically plausible implementation for the type of networks presented in [14].
+Notably, our architecture has the same order of magnitude of trainable parameters and hidden units (see Appendix B).
+In our approach, in-context learning (ICL) can emerge simply by tuning the readout weights of the two involved networks.
+In contrast, previous implementations require biologically implausible mechanisms of temporal credit assignment for the network weights to converge correctly.
+We show that we can side-step this problem by dividing the architecture into two separate components and introducing an additional objective, which forces the first network to approximate the gradient of the second.
+2 Methods
+2.1 Dynamical adaptation replaces synaptic plasticity
 Consider a generic learning task, in which the response $y$ to a state $x$ and is influenced by a set of parameters $w$.
@@
 This can be rewritten in a generic formulation as follows:
-$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f)\\ \tau_{y}\,\dot{y}&=Y(y,w,x)\\ \tau_{w}\,\dot{w}&=W(w,x,e)\end{cases}$ (1)
+$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f),\\ \tau_{y}\,\dot{y}&=Y(y,w,x),\\ \tau_{w}\,\dot{w}&=W(w,x,e).\end{cases}$ (1)
 where $f$ is feedback from the environment (e.g.
@@
 For this reason, we also refer to them as “virtual” weights.
-Dynamical adaptation for linear regression
-Consider now the task of learning a linear relationship between input $x$ and output $y$.
-In the conventional approach, a neural network is initialized with random parameters, essentially making random guesses.
-The model estimation can be improved through iterative adjustments of the network parameters (e.g.
-the synaptic weights), following the gradient of errors evaluated on a training dataset $(x_{i}^{data},y_{i}^{data})$.
-One step of the gradient would update the weights of a linear model $y=wx$, following $\delta w=-\sum_{i}(y-y_{i}^{data})x^{data}_{i}$, where $\eta$ is the learning rate.
-By the end of the training process, the network aims at accurately predict the output $y^{test}$ for a new input $x^{test}$.
-An equivalent error reduction can be achieved by updating the target output $y_{i}^{data}$ to $y_{i}^{data}-wx_{i}^{data}=y_{i}^{data}-y_{i}^{train}$ (let’s denote this the latter quantity as $e_{i}$), where $y_{i}^{train}=wx_{i}^{data}$ is the current model estimation of $y_{i}^{data}$.
-$e_{i}$ can then be updated iteratively, and when it reaches the value zero, it indicates that the interpretation of the dataset is coherent, and the network predictions on new data $x^{test}$ are reliable.
-We can interpret this dynamics as changing the way $x^{test}$ is processed in function of the virtual weights $w$.
-Typically, when a network dynamically alters the way it receives an input, it is referred to as an attention mechanism.
-Following this concept, our goal is to project the dataset $(x_{i}^{data},y_{i}^{data})$ onto the network, influencing the way a new input $x^{test}$ (projected simultaneously) is processed.
-The information that drives and influences the attention mechanisms, $(x_{i}^{data},y_{i}^{data})$ in this case, is usually referred to as a context.
-When the input vector is one dimensional, the task can be solved by a network composed of three modules, one evaluating the error $e$, one evaluating the necessary update $\delta w$, and one evaluating the response $y$ to input data (see Fig.
-1A).
-Their dynamics can be explicitly written as follows  [^1] Here we model the $y$ to be an instantaneous function of the variables $w,x$.
-This is equivalent to considering $Y(y,w,x)=xw-y$ in Equation (1) and taking the limit $\tau_{y}\to 0$.
-Similar considerations hold for $e$.:
-$\begin{cases}e_{i}&=(y^{data}_{i}-x_{i}^{data}w)\\ \tau_{w}\,\dot{w}&=\sum_{i=1}^{M}x_{i}^{data}\,e_{i}\\ y&=x^{test}w\end{cases}$ (2)
-with $e_{i}(0)=y_{i}^{data}-w_{0}x_{i}^{data}$, and $w_{0}$ are the initial weights.
-The conditioning dataset is $(x_{i}^{data},y_{i}^{data})$, with $i=1,...M$, while $(x^{test},y^{test})$ is one element of the test set.
-In this case one example ($M=1$) is sufficient to determine the response to new inputs.
-This formulation is equivalent to what derived in [11], with the difference that here many iteration are recursively evaluated over time, allowing the convergence to the gradient descent solution.
-See Results section for examples of applications.
-This approach can be extended to multidimensional linear datasets (see Supporting Information).
-Dynamical supervised learning for temporal trajectory
-The above formulation is well-posed for stationary dataset, however such a framework (the temporal convergence of the error to zero and of the response to the desired value) is tricky for temporal trajectories, which requires a little change of paradigm, to achieve a formulation that holds at each time step.
-First of all we introduce a target temporal trajectory $y^{targ}(t)$, and define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset.
+2.2 Dynamical supervised learning for temporal trajectory
+We consider the task of learning a target temporal trajectory $y^{targ}(t)$, We define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset.
 In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$.
 We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$.
-This together with Eq. (2) allows us to write:
+In this case, learning can be formulated as follows:
 $\begin{cases}e&=(y^{targ}-y)\\ y&=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{x}\,e\end{cases}$ (2)
@@
 However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
-Dynamical reinforcement learning
+2.3 Dynamical reinforcement learning
 In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments.
@@
 It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
-This formulation holds for a null discount factor [23, 24]).
-See Supplementary information for a complete derivation and for the description of the general case.
+This formulation holds for a null discount factor [23, 24]), we refer to the Appendix section C for the description of the general case.
 The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
-$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (6)
-Algorithm Distillation
+$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}(\bm{y})\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (5)
+2.4 Algorithm Distillation
 To demonstrate the ability of this architecture to learn in context, we train networks Y and W using an Algorithmic Distillation protocol, as defined in [25].
@@
 The network performance is then tested out of distribution, i.e., on the rest of the environments $\mathcal{A}_{\text{OOD}}:=\mathcal{A}\setminus\mathcal{A}_{\text{ID}}$
-Gain modulated reservoir computing (GM-RC)
+2.5 Gain modulated reservoir computing (GM-RC)
 Reservoir Computing
@@
 Consequently, the resulting RNN equation is formulated as:
-$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+R\bm{x})\right)-\bm{z}$ (9)
+$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+\beta\cdot R^{\text{ap}}\bm{x}^{\text{ap}}+R\bm{x})\right)-\bm{z}$ (8)
 Where $\bm{b}^{\text{ap}}$ is a constant bias vector.
+The hyperparameters $\alpha,\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs.
+Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation.
 Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir.
-As before, removing time dependencies we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$.
-We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the effect of the apical inputs on the network.
-Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation.
-The choice for the name $\bm{x}^{\text{ap}}$ is inspired by to current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high level information [22].
-Results
-Figure 1: Dynamical adaptation to changing context: A.
-Network architecture composed of 3 modules: $e$, $\delta w$ and $y$.
-The estimation on new data is produced by $y$, and is modulated by the error variable $e$ through $\delta w$.
-B.
-Example of the time course for the three variables in a single realization of the task.
-The values of $y$ converge to the correct values of (black dashed lines).
-C.
-This model accounts for the recovery of optimal performances (bottom, observed in humans and monkeys data) when the context suddenly changes, compared to a slower increase of performances achieved through synaptic plasticity.
-Dynamical adaptation replaces synaptic plasticity
-To demonstrate that suited responses can be provided relying on dynamical adaptation, we first consider a task requesting to learn a linear relationship between input $x^{data}$ and output $y^{data}$.
-As described in the Methods we consider a three module model (see Fig.
-1A) described by the variables, $e$, $\delta w$, and $y$, whose dynamics is defined by Eq.
-(2), that is equivalent to following a gradient dynamics as respect to the parameters $w_{0}$.
-This update can be realized by treating $e$ as an input that, conditioned by the dataset $(x^{data},y^{data})$, evolves over time.
-When $e\to 0$ the interpretation of the dataset is coherent, and the input-output mapping performed by the network on the new data is reliable.
-In this case one example is sufficient to determine the response to new inputs.
-Fig.
-1B shows the dynamics of three variables: the absolute value of $e$, and $\delta w$ as a consequence the error, are reduced over time, while the $y$ converges to the correct $y^{test}$ given the input $x^{test}$ (shown for 10 different new values).
-We show the performances of the network whose parameters $w_{0}$ are trained through a gradient based plasticity on the rule $y^{data}=x^{data}$ (see first bar in the plots Fig.
-1C), after the rule is switched to $y^{data}=-x^{data}$ (see, the second bar in which performances drop in both panels).
-In the top panel $w_{0}$ parameters are trained through gradient updates (we refer to Github repository https://github.com/cristianocapone/ABSS for details on parameters and implementation).
-In the bottom panel weights are not re-trained, but the last example (that is coherent with the new rule) is projected as a context (e.g.
-$(x^{data},y^{data})=(-1,1)$, starting the trial numbered as 1).
-This represents the agent understanding that the rule has been switched, and allows rapid and dynamic network reconfiguration and recovery of performances to optimal values, similarly to what observed in monkeys and humans data.
-An interesting example is the Wisconsin Card Sorting Test, in which participants are required to sort a deck of cards based on changing rules [30].
-Initially, participants are given a rule (e.g., choosing by color) , and after a certain number of trials, the rule is changed without warning (e.g., choosing by shape).
-Participants then need to adapt to the new rule and adjust their strategy accordingly.
-The quick adaptation observed in this setup [31] reflects the ability to update mental models of the task, and flexibly apply new rules or strategies (similarly to Fig.
-1C-bottom).
-On the other hand in some pathological subject, the adaptation after rule change appears to be slower [31], suggesting a learning process compatible with gradient-based plasticity (Fig.
-1C-top).
+As before, removing time dependencies, we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$.
+We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network.
+The choice for the name $\bm{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [22].
+3 Results
+Computing resources
+The experiments of this paper were executed on a Macbook pro M3 CPU 12-core with 36 GB of RAM, and on a Macbook pro 2,9 GHz 6-Core Intel Core i9 with 32 GB of RAM.
 Figure 1: Dynamical adaptation of temporal trajectories: A.
 Overview of the network architecture employed.
-A recurrent network composed of 10 units (described by the vector $\bm{z}(t)$, depicted in pink) is utilized for learning a periodic trajectory, following the prescription of reservoir computing,.
+A recurrent network composed of $N=20$ units (described by the vector $\bm{z}(t)$, depicted in pink, which follows Eq.
+(6)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing.
 One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights.
@@
 Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics.
-Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights (shown in the green trajectory).
+Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights $\delta w$, a $N$-dimensional vector (shown in the green trajectory).
 Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines).
@@
 C.
-Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios.
+Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (solid: median, dashed: 20-th/80-th percentile (statistics evaluated over 10 realizations).
 Dynamical adaptation for temporal trajectories
-We aim to demonstrate that such an approach can be generalized to non-stationary scenarios.
 We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$.
 Following the reservoir computing paradigm, we employ an RNN (Fig.
-2A, pink), whose dynamics follows Equation (7), to extract temporal features $\bm{z}(t)$.
-During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig.
-2B).
+1A, pink), whose dynamics follows Equation (6), to extract temporal features.
+During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x=y^{targ}(t)$ in Eq.
+(6), while $R$ is Gaussian matrix with zero mean and variance, $\sigma_{R}^{2}$) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig.
+1B).
 This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN.
@@
 These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$.
-We refer to Supporting Information for further details on the training procedure.
-In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN.
+More specifically, the target sequences are $5$ sinusoidal functions with different frequencies (see vertical red lines in Fig.
+1C.)
+In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq.
+(6)).
 This results in an autonomous dynamical system that reproduces the target trajectory.
 (see Fig.
-2B).
+1B).
 In Fig.1B we report an example of successful dynamical adaptation of our model.
@@
 The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.1C and then tested across a range of frequencies.
-Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C).
+Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.1C, solid and dashed lines represent respectively, median and 20th/80th percentile range.
+Statistics is evaluated over 10 realizations of the experiment).
 Figure 2: Dynamical reinforcement learning: Multi Armed Bandits.
@@
 B.
-Comparing in distribution (ID) and out-of-distribution (OOD) expected reward at the 1000-th round for the policy gradient (PG) training source, reward gain-modulated network ($\gamma>0$), and a network without gain modulation $\gamma=0$.
+Regret comparison between the distilled and the original policy gradient algorithm.
+We report log regret per round distribution achieved at the 100-th round for 100 independently trained models in ID and OOD settings.
+We compare a gain-modulated network ($\gamma=1$), and a network without gain modulation ($\gamma=0$) with the policy gradient training source (PG).
 C.
-ID and OOD model performance varying gain modulation strength ($\gamma$).
-Solid lines indicate the median score (expected reward), computed over 20 independently trained models.
-The filled area indicates 25-75% confidence interval.
+OOD model performance varying number of hidden dimensions.
+Solid lines indicate the median score (expected reward), computed over 100 independently trained models.
+The filled area indicates 20-80% confidence interval.
 D.
-Same as B but comparing OOD performance for different network sizes (number of hidden units).
-E.
-Robustness to hyperparameter variations in architectures with $(\gamma=0.9)$ and without ($\gamma=0.0$) gain modulation.
-We report the percentage of models that fail to generalize to OOD environments, varying the standard deviation of apical ($\sigma_{a}$) and basal ($\sigma_{b}$) random input weights.
-F.
-Regret comparison between the distilled and the original policy gradient algorithm.
-We report average regret per round in ID and OOD setting, compared with the one achieved by the PG training source.
-Multi Armed Bandits
+We report average regret per round curves in OOD setting.
+We compare a gain-modulated network ($\gamma=1$, teal), and a network without gain modulation ($\gamma=0$, orchid) with the policy gradient training source (in black).
+Each curve represents the average regret per round for one fixed trained model, computed by averaging over 100 independent simulations.
+3.1 Multi Armed Bandits
 Dynamical adaptation is now investigated within the framework of reinforcement learning.
@@
 We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [30].
-Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability.
+Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability ($p=0.95$).
 During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
@@
 Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig.
-3A, green).
-This serves as an ideal test bed for evaluating the network ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
-To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action, as described in the Methods section.
-We systematically investigate the impact of the gain modulation on out-of-distribution performance, varying the parameter $\gamma$.
-Notably, when $\gamma=0$ the reward does not modulate the network gain.
-We present our experimental findings in Fig.
-3.
-Initially, we observe that integrating reward gain modulation into the network enables the agent to consistently converge to the optimal policy, even in out-of-distribution environments (see Fig.3B).
-To further explore this phenomenon, we trained multiple models while varying the strength of the modulation parameter ($\gamma$).
-We observe that performance improves with higher values of $\gamma$, allowing the generalization of learned rules out-of-distribution (Fig.
-3C) and showcasing increased robustness to variations in hyperparameters (Fig.
-3E).
-Furthermore, we observe that in the absence of gain modulation, the network is unable to generalize, merely memorizing the rules from the training set.
-This is evidenced by its persistence in selecting even-numbered arms even when tested in out-of-distribution environments where high rewards are allocated to odd-numbered arms (see Fig.
-3E).
-Comparing the performances of models of different sizes (number $N$ of hidden units), we see that models with more reward gain modulation need significantly fewer neurons to achieve maximum scores in OOD environments (see Fig.3D).
-Analysing the regret curves displayed in Fig.
-3F we observe that a model with gain modulation learns a more data-efficient algorithm than its source even in OOD environments.
-On the contrary, a model without gain modulation fails to generalize to OOD environments, not converging to zero regret (see Fig.
-3F, right panel).
+2A, green).
+This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
+To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action.
+We train this network to approximate the policy gradient update rule.
+As training data, we use policy gradient estimates over one single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment.
+We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) models in out-of-distribution (OOD) environments.
+We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment.
+We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) in out-of-distribution (OOD) environments.
+To systematically investigate the impact of gain modulation ($\gamma=1$) on out-of-distribution performance, we compare it with networks where the reward does not modulate the network gain ($\gamma=0$).
+For each case, we select the optimal hyperparameters through a grid-based search (additional details in the Appendix).
+Our experimental findings are presented in Fig.2.
+First, comparing models of different sizes (number $N$ of hidden units), we find that models with gain modulation require significantly fewer neurons to achieve maximum scores in OOD environments (Fig.2C).
+In Fig.2B, we report the regret per round distribution at the 100th round and compare it with the distribution obtained by policy gradient with the same learning rate and iterations.
+A gain-modulated network achieves a regret distribution comparable to the policy gradient in both ID and OOD environments, with a lower median.
+In contrast, networks without gain modulation show significantly higher regret.
+Analyzing the regret curves in Fig.2D, we observe that a model with gain modulation often learns a more data-efficient algorithm than its source, even in OOD environments.
+Conversely, a model without gain modulation fails to generalize to OOD environments and does not converge to zero regret.
+In summary, gain modulation enables the network to consistently and efficiently distill the correct gradient update rule and generalize it to unseen environments, predicting the correct virtual weight update in regions of the input space far from the training data.
 Figure 3: Dynamical Reinforcement Learning Dark Room A.
@@
 C.
-Illustration of the task: the agent begins from the center and learns to reach unobservable objects positioned at various locations (represented by colored circles).
+Illustration of the task: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles).
 After several trials, the agent achieves precise targeting of the circles.
 Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel).
-E.
-Reward plotted against the number of trials, for the policy gradient (PG) algorithm (displayed in blue lines) applied in the dynamical PG scenario on both training positions (depicted in the left panel) and new positions (shown in the right panel).
-Thin line: single position, thick line: average across different positions.
+C.
+Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD).
+Line: median, shading: 20-th/80-th percentile range.
 B, D, F.
 Similar to A, C, E, respectively, but without gain modulation.
-Reinforcement learning in a reaching task.
+3.2 Reinforcement learning in a reaching task.
 We examine a reaching task, known as the dark room task (a simple instance of the water maze task [31]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$.
@@
 Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility.
-The horizontal and vertical positions of the agent are encoded separately using five input units $\bm{x}$ each, employing Gaussian activation functions with equally spaced centers ($0.4$ apart) and a width of $0.4$.
-To achieve this task we consider a network architecture composed of two networks.
-One GM-network, , $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig.
-4A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig.
-4A, shown in orange) resulting from changes in virtual weights.
+The position of the agent is encoded separately using 25 input units $\bm{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
+To accomplish this task we consider a network architecture composed of two networks.
+One GM-network, $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig.
+3A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig.
+3A, shown in orange) resulting from changes in virtual weights.
 Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section.
-Their dynamics can be described by the following equations:
-$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq\bm{x}\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\end{cases}$ (11)
-Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
+Their dynamics can be described by the following equation Eq.(5).
+Indeed, we compare the case with (Fig.
+3A) and without (Fig.
+3B) gain modulation.
+We refer to Supporting Information for further details on the training procedure.
 Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig.
-4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
-4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
+3C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
+3E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
 After multiple trials, the agent successfully achieved precise targeting of the circles.
@@
 To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig.
-4C, left panel) and new test positions (OOD, Fig.
-4C, right panel).
+3C, left panel) and new test positions (OOD, Fig.
+3C, right panel).
 For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training.
@@
 We present the reward plotted against the number of trials for both training (Fig.
-4E, right panel, pink lines) and test (Fig.
-4C, right panel, pink lines) food locations.
+3E, right panel, pink line) and test (Fig.
+3C, right panel, pink line) food locations.
 We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig.
-4B, D, F).
+3B, D, F).
 Notably, while performances for ID food locations are acceptable (Fig.
-4E, left panel, pink lines), those for OOD cases are extremely low (Fig.
-4E, right panel, pink lines).
+3E, left panel, pink lines), those for OOD cases are extremely poor (Fig.
+3E, right panel, pink lines).
 This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
-Discussion
+We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section C in Appendix), requiring evaluating temporal credit assignment.
+4 Discussion
 It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 32], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8].
 Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers.
-One intriguing explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11].
+This explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11].
 However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals.
-In this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
-When considering a stationary dataset our model dynamically adapt to the context provided by the data, allowing predictions on new data, without synaptic plasticity.
-Remarkably, when the learning rule shifts abruptly, the network swiftly adapts without retraining, mimicking human cognitive flexibility observed in tasks like the Wisconsin Card Sorting Test.
-Conversely, when using gradient-based plasticity, the learning speed resembles what observed in certain pathological conditions [31].
+In this this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
 When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity.
@@
 Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
-In conclusion, our approach provide an explicit framework to induce ICL in biologically plausible network, possibly opening the route to a formal understanding of ICL in biological agents.
-Acknowledgments
+In conclusion, our approach provides an explicit framework to induce ICL in biologically plausible networks, possibly opening the route to a formal understanding of ICL in biological agents.
+Limitations of this study
+In this study, we focused on dynamical adaptation of virtual readout weights, in other words, only the last layer is considered.
+We are neglecting the dynamics on recurrent weights and weights of the hidden units, which, in general, strongly affect performance and stability.
+Generalizability to untested real-world conditions is uncertain, and we do not investigate the scalability to larger architectures.
+We discuss that the mechanism we discuss could be present in biological agents, allowing in-context learning.
+However, the theoretical basis for attention mechanisms executing gradient descent updates lacks comprehensive empirical validation in biological experiments.
+Additionally, the study does not fully address how dynamic gradient structures could be learned or developed in biological networks.
+5 Acknowledgments
 This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’) to MM.
@@
 The source code is available under CC-BY license in the https://github.com/cristianocapone/ABSS public repository.
+Appendix A Relationship with previous work on product-based architectures implementing in context-learning
+The relationship between attention-based transformer architectures and in-context learning was first noted in [11] where it was shown through constructive proof, confirmed by experiments, that linear attention layers can implement a gradient descent update for linear regression in its forward pass.
+Building upon this insight, [14] showed that a linear gated RNN can implement the same mechanism, showing that a linear two-layer gated RNN can replicate a linear transformer.
+The work proposes an implementation that uses $O(d^{2})$ hidden units and has $O(d^{4})$ trainable weights, that can be reduced to $O(d^{3})$ by side gating.
+Our gain modulated architecture significant analogies with the gated RNNs analyzed in [14].
+Consider a linear  [^3] For simplicity, here we assume $\alpha,\beta=0,\ \gamma=1$ and consider a linear activation function $\phi=Id$.
+gain-modulated network architectures implementing the functions ${\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})$, ${\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)$, dynamically performing a ICL linear regression task that requires to map temporal dependent features $z(t)\in\mathbb{R}^{N_{in}}$ to predict an output $\bm{y}^{targ}\in\mathbb{R}^{N_{out}}$:
+$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})=\Theta^{y}\left(R^{ap}_{y}\bm{w}\right)\odot(R_{y}\bm{z}(t))\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},\bm{y}^{targ}-\bm{y})=\Theta^{w}\left(R^{ap}_{w}\left(\hat{y}-y\right)\right)\odot(R_{w}\bm{z}(t))\end{cases}$ (10)
+This architecture can be seen as a particular instance of the RNN with side gating (see Equation (15) in [14]).
+Interestingly, a similar mechanism is also present in the recently proposed Mamba layer [33].
+Appendix B Approximating gradients with gain modulated architectures
+In this appendix, we show that an architecture with gain modulation is more suited to approximate gradient terms involved in dynamical learning.
+Scalar product
+We first consider the task of approximating a scalar product between virtual weights $\bm{w}\in\mathbb{R}^{N_{in}}$ and features $\bm{x}\in\mathbb{R}^{N_{in}}$.
+To achieve this, we train the readout weights $\Theta$ of a gain modulated network $\mathrm{GMNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{w}\})$ with $N_{h}$ hidden features to approximate the function $dot(\bm{x},\bm{w})=\sum_{i=1}^{N_{in}}x_{i}w_{i}$.
+The training dataset is composed of $1000$ pairs $(\bm{x},\bm{w})$ uniformly sampled in the hypercube $[0,1]^{2N_{in}}$.
+The test set consists of the same number of pairs sampled in the hypercube $[-1,1]^{2N_{in}}$.
+We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
+For each of the two settings, we select the best hyperparameters by fixing the dimensionality of the inputs ($N_{in}=5$) and hidden units ($N_{h}=101$) and varying the standard deviation of the projection matrices $R_{ij},R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}),\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
+The best hyperparameters found were used in the models to test the approximation performance of the dot product varying the number of input features ($N_{in}\in[1,10]\cap\mathbb{N}$) and hidden units ($N_{h}\in[1,200]\cap\mathbb{N}$).
+Results are shown in Fig.
+A1 A, B, C.
+We see that while both architectures require a quadratic number of hidden units to achieve low error on the test set, the gain-modulated network after a threshold number of units is hidden units is reached, is able to perfectly approximate the scalar product function consistently achieving $\sim 10^{-7}$ RMSE error.
+At the same time, the error in the network without gain modulation remains several orders of magnitude higher.
+Scalar-vector product
+We then tackle the task of approximating a vector-scalar product between features $\bm{x}\in\mathbb{R}^{N_{in}}$ and features and error $\bm{e}\in\mathbb{R}$.
+To achieve this, we train the readout weights $\Theta$ of a gain modulated network $\mathrm{GMNN}_{\Theta}(\{\bm{x}\}|\{\gamma e\})$ with $N_{h}$ hidden features to approximate the function $prod(\bm{x},e)=e\bm{x}\in\mathbb{R}^{N_{in}}$.
+The training dataset is composed of $1000$ pairs $(\bm{x},e)$ uniformly sampled in the hypercube $[0,1]^{N_{in}}\times[0,1]$.
+The test set consists of the same number of pairs sampled in the hypercube $[-1,1]^{N_{in}}\times[-1,1]$.
+We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
+For each of the two settings, we select the best hyperparameters by fixing the dimensionality of the inputs ($N_{in}=5$) and hidden units ($N_{h}=21$) and varying the standard deviation of the projection matrices $R_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}/N_{in}),R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2},\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
+The best hyperparameters found were used in the models to test the approximation performance of the dot product varying the number of input features ($N_{in}\in[1,10]\cap\mathbb{N}$) and hidden units ($N_{h}\in[1,40]\cap\mathbb{N}$).
+Results are shown in Fig.
+A1 D, E, F.
+We see that the network without gain modulation requires a quadratic number of hidden units to achieve low error on the test set.
+As in the scalar product case, the gain-modulated network can nearly perfectly approximate the target function after a threshold number of hidden units is reached.
+Significantly, in this case, the threshold scales linearly with $\mathbb{R}^{N_{in}}$.
+Figure A1: A Test errors for the scalar product approximation task, varying the number of features $N_{in}$, hidden units $N_{h}$.
+We compare architectures with gain modulation ($\gamma=1$, bottom) with architectures without gain modulation ($\gamma=0$, bottom).
+B Test errors for the dot product approximation task, varying the number of features $N_{in}$, and fixing the number of hidden units $N_{h}=101$.
+Solid lines indicate the median over 100 trained models, while the filled region indicates the $20/80th$-percentile interval.
+C Same as B but fixing the number of features $N_{in}=5$, and varying the number of hidden units $N_{h}=101$.
+D, E, F Same as panels A, B, C but for the scalar-vector product approximation task.
+On the number of hidden units and trainable weights needed in our model
+Consider the architecture in Eq. (10) with $N_{out}=1$:
+$\begin{cases}y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})=\Theta^{y}\left(R^{ap}_{y}\bm{w}\right)\odot(R_{y}\bm{z}(t))\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)=\Theta^{w}\left(R^{ap}_{w}\left(\hat{y}-y\right)\right)\odot(R_{w}\bm{z}(t))\end{cases}$ (11)
+Where $R^{ap}_{y},R_{y},R^{ap}_{w},R_{w}$ are fixed random matrices respectively of dimension $N_{y}\times N_{in},N_{y}\times N_{in},N_{w}\times 1,N_{w}\times N_{in}$, and $\Theta^{y},\Theta^{w}$ are trainable readout matrices of sizes $1\times N_{y},N_{in}\times N_{w}$.
+In this simplified linear setting, the features extracted by the network are linear combinations of the product of the input entries that need to be linearly combined by the readout weights to the target function, which in both cases is composed of a specific linear combination of these products.
+It is then straightforward to observe that, to approximate the dot product, in ${\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})$ we need $N_{w}\sim O(N_{in}^{2})$ hidden units and $O(N_{in}^{2})$ readout parameters.
+as confirmed in the experiment.
+This number can be further reduced to $O(N_{in})$ both for the hidden units and trainable weights  [^4] This can be done assuming that $R^{ap},R_{w}$ are diagonal matrices (and thus reducing the problem to estimating $N_{in}$ $1\times 1$ products).
+As for the network ${\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)$ approximating vector-scalar products, similar considerations support the experimental observation that $N_{y}\sim O(N_{in})$ hidden units and $O(N_{in}^{2})$ features are needed in this case.
+For a multivariate $\mathbb{R}^{N_{in}}\to{N_{out}}$ linear regression task, we can consider $N_{out}$ independent modules of the type described before.
+In this case, the number of trainable parameters is $O(N^{2}_{in}\cdot N_{out})$ with $O(N_{in}\cdot N_{out})$ hidden units ($O(N_{in}^{2}\cdot N_{out})$ in the more biologically plausible case).
+Appendix C Reinforcement Learning and the Discount Factor
+Reinforcement Learning (RL) is a machine learning paradigm where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards.
+The discount factor, usually denoted by $\gamma$ (where $0\leq\gamma\leq 1$), is crucial in RL as it determines the importance of future rewards.
+The cumulative reward $R_{t}$ at time step $t$ is given by:
+$R_{t}=r_{t}+\gamma r_{t+1}+\gamma^{2}r_{t+2}+\gamma^{3}r_{t+3}+\ldots$
+In our experiment, the policy gradient update rule in the presence of the discount factor could be approximated as (see [23]):
+$\begin{cases}\bm{e}&=\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=r\bm{\hat{e}}\end{cases}$ (12)
+where $\bm{\hat{e}}$ is an exponential temporal filtering of $\bm{e}$, with a timescale that is proportional to $-\frac{1}{\log(\gamma)}$.
+The importance of the discount factor lies in its ability to balance immediate and future rewards.
+A higher $\gamma$ values future rewards more significantly, encouraging the agent to consider the long-term consequences of its actions.
+Conversely, a lower $\gamma$ makes the agent prioritize immediate rewards.
+This balance is essential for the stability of the learning process.
+An appropriately chosen $\gamma$ ensures stable learning and convergence; if $\gamma$ is too high, the agent might overvalue distant future rewards, leading to instability, while a very low $\gamma$ can result in short-sighted behavior.
+If the discount factor is zero, the agent’s policy would change only when it is very close to the moment the reward is received.
+In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points.
+To demonstrate that our approach performs the temporal computation required to consider future rewards, we compare the performances of our network (see Fig.
+A2.A-B) to a network without recurrent connections, which therefore cannot perform temporal computations.
+Recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions.
+Without these connections, a network operates purely on the current input without any contextual information from prior steps, thus lacking the ability to perform temporal computations (see Fig.
+A2.D-E).
+Performances are higher in the first case.
+This can be visualized by looking at the policy at the end of the dynamic reinforcement learning for a specific target location.
+When the recurrent weights are set to zero, the policy points towards the target position only when nearby the target itself (see Fig.
+A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig.
+A2F, black line).
+On the other hand, in the presence of recurrent weights, the proper policy (pointing towards the target, see Fig.
+A2C) is known even when far from the target, allowing optimal long-term planning (see Fig.
+A2C, black line).
+Figure A2: Discount Factor and eligibility traces: A.
+Sample trajectories of the agent at the end of the learning procedure, for various target locations (color-coded), different from the ones used to train the gradient and the scalar product networks (red crosses).
+B.
+Reward as a function of number of trials (or games), blue: policy gradient, pink: dynamical learning for new positions (OOD).
+Line: median, shading: 20-th/80-th percentile range.
+C.
+Arrows: policy at the end of the training as a function of the position.
+Line: sample trajectory of the agent at the end of learning.
+Small circle: agent position, double circle: target position.
+D-E-F same as in A-B-C, but the recurrent connections of the gradient network are set to zero.
+Appendix D Additional details on the "dynamical learning of a temporal trajectory" experiment
+In the "dynamical learning of a temporal trajectory" experiment we test our architecture and non-synaptic learning approach on temporal tasks.
+The primary goal is to analyze the network’s ability to generalize beyond the target frequencies used to pre-train our networks.
+Table 1: Simulation Parameters
+Parameter Symbol Reservoir Gradient Net Scalar Net
+Network Size $N$ 20 500 500
+Input Dimension $I$ 1 100 + 10 100+100
+Apical Input Dimension $I^{ap}$ 0 1 100
+Output Dimension $O$ 1 100 1
+Time Step $dt$ 0.005
+Reservoir Time Constant $\tau_{m_{f}}$ 10 $dt$ 1 $dt$ 1 $dt$
+Input weights var $\sigma_{input}$ 0.06 0.06 0.06
+Apical Input weights var $\sigma_{input}^{ap}$ 0.0 0.1 0.1
+Recurrent weights $\sigma_{rec}$ 0.99 / $\sqrt{N}$ 0.5 / $\sqrt{N}$ 0. / $\sqrt{N}$
+Gain-modulation factor $\gamma_{net}$ 0. 1. 1.
+Three networks are defined in this experiment: one uses a reservoir to compute temporal features and encode the target temporal trajectory, one predicts the required weight updates, and one predicts the scalar product, as discussed in the main text.
+The parameters used in the simulation are summarized in Table 1.
+Inputs are projected to the network through Gaussian weights with zero mean and variance $\sigma_{in}^{2}$ to distribute the input information across multiple units in the reservoir.
+Data collection and pre-training
+The readout of the first reservoir network is defined as $R_{j}$, and the readout at time $t$ is given by:
+$y(t)=\sum_{j}R_{j}z_{j}(t),$
+where $z_{j}(t)$ are the reservoir states.
+Training and evaluation involve presenting the networks with five target trajectories $y^{targ}(t)=0.8sin(\omega_{targ}t)$, with five angular velocities, $\omega_{targ}$, ranging from $0.04$ to $0.08$.
+The reservoir readout parameters are trained using online gradient descent to minimize the error between the current output and the target output.
+The loss function is defined as:
+$L=\frac{1}{2}\left(y(t)-y^{\text{target}}(t)\right)^{2}.$
+The gradient used for online training is:
+$\Delta R_{j}=-\eta\frac{\partial L}{\partial R_{j}}=-\eta\left(y(t)-y^{\text{target}}(t)\right)z_{j}(t),$
+where $\eta$ is the learning rate.
+During data collection, trajectories of network states, errors, and weight updates are recorded.
+The gradient and scalar-product networks are trained on this data to predict the required gradient given a specific temporal error, and to estimate the current trajectory given the current virtual weights.
+The networks are pre-trained by estimating only their readout weights, in accordance with a reservoir computing prescription. They are trained using a pseudo-inverse approach.
+Dynamical supervised online learning and evaluation
+Once the architecture is pre-trained, it can be tested to learn new trajectory, unobserved during the pretraining.
+readout weights of the reservoir are no longer changed online using gradient, but rather virtual weights are changed following the prescription of the gradient network.
+Virtual weights are then used as an input to the scalar-product network to predict the current prediction $y(t)$.
+Performances are evaluated by measuring the MSE between the target trajectory ant the predicted one $y(t)$, for different values of the trajectory angular velocity, equally distribute between 0.01.01.01.01 and $0.1$.
+Appendix E Additional details on the bandit experiment
+Family of tasks
+We consider a family of tasks $\mathcal{A}$, such that every element $\alpha\in\mathcal{A}$ represents a Bernoulli $K$-armed bandit problem with rewards $\{R_{i}^{\alpha}\}_{i=1}^{K}$.
+Each task $\alpha$ is specified by a set $P_{\alpha}\subset\{1\cdot,K\}$ of positive arms, such that
+$R_{i}^{\alpha}=\begin{cases}Ber({p})&i\in{P_{\alpha}}\\ Ber({1-p})&i\in{P_{\alpha}}\end{cases}$ (13)
+In the main text experiment, we used $K=10$ and $P_{ID}:=\{i\in[K]|i\equiv 0(\text{mod}2)\}$ and $P_{OOD}:=\{i\in[K]|i\equiv 1(\text{mod}2)\}$.
+Network details and hyperparameter search
+We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
+For each of the two settings, we select the best hyperparameters by fixing the number of hidden units ($N_{h}=100$) and varying the standard deviation of the projection matrices $R_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}/20)\ \ R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}),\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
+For each point in the grid, we train 20 models using The best hyperparameters found were used in the subsequent experiments In Fig.
+2 B, D the reported regrets are on models trained with $N_{h}=100$ hidden features
+Regret per round
+Given an agent that plays in an environment $\alpha$ receiving rewards $\{r_{t}\}_{t\in\mathbb{N}_{+}}$, the regret per round $\rho(T)$ achieved by the agent at round $T$ is defined as:
+$\rho(T)=\mu^{\alpha}_{\star}-\frac{1}{T}\sum_{t=1}^{T}r_{t}$ (14)
+Where $\mu^{\alpha}_{*}:=\max\{\mathbb{E}[R_{k}^{\alpha}]|k\in[K]\}$ is the maximum expected reward that can be obtained in each round playing the optimal policy that selects one of the best arms with probability $1$.
+Appendix F Additional details on the darkroom experiment
+This experiment investigates the capability of out dynamical reinforcement learning approach, to improve an agent’s ability to navigate a 2D environment.
+The agent is trained to locate a randomly placed, not observable food source, with its policy evolving over multiple episodes through.
+The environment is a 2D grid where an agent starts at the center, aiming to reach a randomly positioned food item.
+The food positions vary every 600 episodes, introducing.
+The agent’s movements are restricted within the grid, with actions limited to moving left, right, up, or down.
+To represent the agent’s position, a place cell encoder converts the agent’s $(x,y)$ coordinates into a higher-dimensional feature vector using Gaussian functions.
+This encoding helps in effectively capturing spatial information.
+The agent’s policy is linear, with action probabilities derived from a softmax function applied to the encoded state.
+For this reason, only two networks are used for the task, the gradient and the scalar-product network see Table 2 for details on parameters.
+Data collection and pre-training
+The policy is updated using the policy gradient method, adjusting weights based on received rewards to improve decision-making over time.
+Learning data are collected for 8 different positions of the food, equally distributed on a circle.
+Two types of networks are used: one to model the agent’s state dynamics and another to handle gradient updates necessary for learning.
+These networks influence the agent’s internal state and learning process, enabling policy refinement.
+The two networks readout are trained with a linear regression on their readout weight, to reproduce the proper weight updates an scalar products.
+Dynamical reinforcement learning and evaluation
+The agent is tested again on the 8 positions used for pre-training, and on 8 new positions.
+Similarly to previous experiments, policy gradient is no longer used and replaced by virtual weights updated estimated by the gradient network, and used by the scalar-product network to predict agent policy.
+Performance is measured by the total reward accumulated over episodes. Visualizing the agent’s trajectories reveals its navigation efficiency and decision-making process.
+Table 2: Simulation Parameters
+Parameter Symbol Gradient Net Scalar Net
+Network Size $N$ 500 1000
+Input Dimension $I$ 25+4+10 4 $\times$ 25+4 $\times$ 25
+Apical Input Dimension $I^{ap}$ 1 4 $\times$ 25
+Output Dimension $O$ 4 $\times$ 25 4
+Time Step $dt$ 0.005
+Reservoir Time Constant $\tau_{m_{f}}$ 1 $dt$ 1 $dt$
+Input weights var $\sigma_{input}$ 0.01 0.01
+Apical Input weights var $\sigma_{input}$ 0.1 0.1
+Recurrent weights $\sigma_{rec}$ 0.5 / $\sqrt{N}$ 0. / $\sqrt{N}$
+Gain-modulation factor $\gamma_{net}$ 1. 1.

### v4

@@
 Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid.
-Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters.
-However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration.
+Improvement in behavioral performances are usually associated learning theories to synaptic plasticity.
+However, such rapid changes are not coherent with the timescales of biological synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical reconfiguration of the network involved, without changing its weights.
 In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing.
 Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters.
-Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks.
-We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks.
+Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks, such as in pyramidal neurons thanks to input segregation and dendritic amplification.
+We propose a constructive approach to induce in-context learning in an architecture composed of recurrent networks with gain modulation, demonstrating abilities inaccessible to standard networks.
+In particular, we show that, such architecture can dynamically implement standard gradient-based by encoding weight changes in the activity of another network.
+We argue that, while these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the level of neural circuits.
 We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning.
 Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
-1 Introduction
-The study of behavioral adaptation, observed in both humans and animals, has been a longstanding subject of research.
-The rapidity with which behaviors can change in response to new cues challenges the conventional explanation of synaptic plasticity.
-This led researchers to explore mechanisms that underlie such flexible adaptations [1].
+Introduction
+Adaptive Behavior in Biological Agents
+The study of adaptive behavior in biological agents, including humans and animals, has been a longstanding subject of research [1].
+Observations of rapid behavioral shifts in response to environmental cues have raised questions about whether traditional mechanisms, such as synaptic plasticity, can fully account for such flexibility.
+A broad range of neural and behavioral data indicates the existence of distinct systems for behavioral choice [2, 3].
+The conventional view posits that subcortical areas support habitual, stimulus-based responses, enabling reflex-based and rapid reactions.
+In contrast, higher cortical areas, such as the prefrontal cortex, are associated with reflective, goal-directed behaviour, which relies on cognitive action planning and deliberate evaluation.
+Critically, goal-directed behavior allows for the flexible construction of value at each decision point through online planning procedures.
+This mechanism makes agents immediately sensitive to changes in outcome values, enabling them to dynamically adapt their actions to novel or changing circumstances [4].
+Studies in network neuroscience have identified dynamic network reconfiguration, characterized by the flexible recruitment and integration of neural circuits, as a fundamental mechanism supporting this adaptability across cognitive[5] and motor domains[6].
+Computational models within the predictive coding framework have recently provided insights into the neural mechanisms underlying dynamic adaptation.
+These models suggest that adaptation can be understood as a form of Bayesian inference, where the brain continuously infers latent contextual states to guide behavior [7].
+A prominent example is the COIN model[8, 9], which posits that adaptation arises through two complementary processes: proper learning, involving the creation and updating of memories, and apparent learning, which dynamically adjusts the expression of pre-existing memories based on inferred context.
+While these frameworks capture key phenomena of adaptive behavior, such as context-dependent single-trial learning and spontaneous recovery, a significant gap remains between these theoretical models and the neural substrates that implement these computations in the brain.
 In-context learning as an emerging property in AI
-In-context learning (ICL) is the capacity of a model to adapt its behavior, without weight updates, to solve tasks not encountered during training.
+In machine learning, the capacity of a model to adapt its behavior, without weight updates, is referred to as in-context learning (ICL).
 Initially, ICL was observed in architectures tailored for few-shot learning [10] or even zero-shot learning [11].
 However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [12, 13].
+These architectures leverage self-attention, a mechanism that dynamically changes the way past inputs are attended, in order to shape the current response.
 Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[14, 15, 16].
-Despite some intriguing explanations [9, 10, 11], the emergence of this property remains not completely understood.
-In particular, it is not clear how to associate it with in-context learning abilities observed in biological networks.
-The major contribution of this work is the definition of a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
-Limitations of transformers architecture
-Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length.
-This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding.
-To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs.
-Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks.
-Zucchet et al.
-[14] explore the efficacy of these deep linear gated RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and implementing in-context supervised learning.
+While there are several intriguing explanations [17, 18] for the emergence of ICL in transformers, a more formal understanding was pioneered by Von Oswald and collaborators [19], who demonstrated that a linear self-attention layer can implement gradient descent steps within its forward pass.
+This intuition was empirically validated by analyzing trained transformers and was formally proven [20].
+Despite their success, transformer architectures face notable shortcomings, particularly in their memory requirements, which scale quadratically with sequence length.
+These requirements make those architectures largely biologically implausible.
+Moreover, memory demands constrain their scalability and restrict their applicability to tasks involving long sequences, such as full-length document analysis or video understanding.
+To overcome these challenges, several efficient deep linear RNN architectures have been proposed [21, 22], demonstrating substantial performance gains over transformers, especially in long-sequence tasks.
+However, these deep learning architectures rely on several mechanisms that do not have an obvious correspondence in biological neural systems (such as the attention mechanism itself), leaving an unresolved gap in relating them to the in-context learning abilities exhibited by biological networks.
+This work bridges this gap by introducing a constructive method to induce in-context learning in biologically plausible neural networks, demonstrating its applicability in a wide range of scenarios.
 Biological support for in-context learning
-Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks.
+Building on this premise, our investigation explores how in-context learning, as demonstrated in deep learning architectures, might also be realized in biological recurrent neural networks.
 Are there unique features in transformers that are also present in biological networks?
-We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks.
-We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers.
+We propose that in-context learning can be facilitated by input segregation and dendritic amplification, two features widely observed in biological networks.
+We further argue that these mechanisms provide a biologically plausible foundation for implementing a process analogous to the attention mechanism in transformers.
 Recent findings on dendritic computational properties [23] and on the complexity of pyramidal neurons dynamics [24] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [25, 26, 27, 28].
-It has has been proposed that segregation of dendritic input [18] (i.
+It has been proposed that segregation of dendritic input [26] (i.
 e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [28] would support backpropagation in biological neurons.
-In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks.
+Recently, it has been suggested [29] that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks.
 They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network.
@@
 We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations.
-This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behaviour in both natural and artificial intelligence domains.
-Our work generalizes and provides a biologically plausible implementation for the type of networks presented in [14].
-Notably, our architecture has the same order of magnitude of trainable parameters and hidden units (see Appendix B).
-In our approach, in-context learning (ICL) can emerge simply by tuning the readout weights of the two involved networks.
-In contrast, previous implementations require biologically implausible mechanisms of temporal credit assignment for the network weights to converge correctly.
-We show that we can side-step this problem by dividing the architecture into two separate components and introducing an additional objective, which forces the first network to approximate the gradient of the second.
-2 Methods
-2.1 Dynamical adaptation replaces synaptic plasticity
-Consider a generic learning task, in which the response $y$ to a state $x$ and is influenced by a set of parameters $w$.
-The latter are adjusted in function of an internal error $e$.
-This can be rewritten in a generic formulation as follows:
-$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f),\\ \tau_{y}\,\dot{y}&=Y(y,w,x),\\ \tau_{w}\,\dot{w}&=W(w,x,e).\end{cases}$ (1)
-where $f$ is feedback from the environment (e.g.
-the reward, the target behavior, …).
-Alternatively to the standard interpretation of $w$ as synaptic weights of a neuronal network, we consider them as the state variables of a dynamical system coupled to the dynamics of $y$, expressed by the activity of an auxiliary network.
-For this reason, we also refer to them as “virtual” weights.
-2.2 Dynamical supervised learning for temporal trajectory
-We consider the task of learning a target temporal trajectory $y^{targ}(t)$, We define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset.
-In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$.
-We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$.
-In this case, learning can be formulated as follows:
-$\begin{cases}e&=(y^{targ}-y)\\ y&=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{x}\,e\end{cases}$ (2)
-where we removed the dependence on time $t$ for simplicity.
-The operations required are nonlinear, and usually are naturally implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights.
-However, here $\bm{w}$ are not actual weights but rather dynamical variables.
-We propose, as possible implementation of this, that such non-linear functions are computed by two neural networks networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
-$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{x},e)\simeq W(\bm{x},e)=\bm{x}\,e\end{cases}$ (3)
-In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm).
-In addition, we introduce a novel concept known as a gain-modulated network.
-This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]).
-These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions.
-More details in the section "Gain Modulated Reservoir Computing".
-We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents.
-The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$.
-However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
-2.3 Dynamical reinforcement learning
-In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments.
-We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions.
-For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$.
-This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above.
-The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN.
-This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time.
-This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
-$\begin{cases}\bm{e}&=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{e}\odot\bm{x}\end{cases}$ (4)
-This is the dynamics obtained by evaluating the policy gradient with respect to the virtual weights $\bm{w}$.
-$a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones.
-$\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [23, 24]) at time $t$.
-It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
-This formulation holds for a null discount factor [23, 24]), we refer to the Appendix section C for the description of the general case.
-The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
-$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}(\bm{y})\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (5)
-2.4 Algorithm Distillation
-To demonstrate the ability of this architecture to learn in context, we train networks Y and W using an Algorithmic Distillation protocol, as defined in [25].
-Namely, we consider a family of environments $\mathcal{A}$ with different reward distributions  [^1] A family of environments is a set of possible environments with the same action-state transitions but different reward distributions..
-We then train the aforementioned networks using $N_{s}$ learning histories obtained through policy gradient methods from a subset of these environments, designated as $\mathcal{A}_{\text{ID}}$.
-The network performance is then tested out of distribution, i.e., on the rest of the environments $\mathcal{A}_{\text{OOD}}:=\mathcal{A}\setminus\mathcal{A}_{\text{ID}}$
-2.5 Gain modulated reservoir computing (GM-RC)
-Reservoir Computing
-Reservoir Computing (RC) represents a paradigm in machine learning that provides a model for biologically plausible computation and learning, drawing inspiration from the information processing mechanisms observed in biological neural networks.
-In this approach, a random RNN is employed to extract features from a time-dependent signal $\bm{x}(t)$.
-The dynamics of the RNN describe the evolution of $N$ hidden units $\bm{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
-$\tau_{z}\dot{\bm{z}}=\phi\left(J\bm{z}+R\bm{x}\right)-\bm{z}$ (6)
-The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([26]).
-Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively.
-Subsequently, the features extracted by the network can be utilized to predict a target signal $\bm{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\bm{y}^{targ}(t)-\Theta\bm{z}(t)|^{2}_{2}$ is minimized.
-The reservoir is then implementing a map
-$\bm{y}(t)=\mathrm{RNN}_{\Theta}(\{\bm{x}(s)\}_{s\leq t})$ (7)
-where $\Theta$ represents trainable parameters.
-In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies.
-222 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
-When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit.
-In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\bm{x}\right)$.
-Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights.
-This architecture is also referred to in the literature as the Extreme Learning Machine [28].
-Gain modulated network architecture
-Analysing Equations (3) and (5) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters.
-Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC).
-This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [22, 29].
-Here, we consider an additional input source $\bm{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$.
-Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope.
-Consequently, the resulting RNN equation is formulated as:
-$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+\beta\cdot R^{\text{ap}}\bm{x}^{\text{ap}}+R\bm{x})\right)-\bm{z}$ (8)
-Where $\bm{b}^{\text{ap}}$ is a constant bias vector.
-The hyperparameters $\alpha,\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs.
-Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation.
-Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir.
-As before, removing time dependencies, we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$.
-We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network.
-The choice for the name $\bm{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [22].
-3 Results
-Computing resources
-The experiments of this paper were executed on a Macbook pro M3 CPU 12-core with 36 GB of RAM, and on a Macbook pro 2,9 GHz 6-Core Intel Core i9 with 32 GB of RAM.
+This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
+Main / Framework description
+Dynamical adaptation replaces synaptic plasticity: A popular example
+Consider a task inspired by the Wisconsin card-sorting test.
+Each card in a deck is defined by two features: color and shape.
+In each iteration, the agent receives a cue card and two guide cards—one matching the cue by color and the other by shape.
+The agent must select the correct guide card to receive a positive reward $R(t)=1$.
+Initially, the correct choice is based on color matching.
+However, after 10 trials, the classification rule changes, rewarding shape matching instead.
+A successful agent must quickly adapt its decision-making to these contextual shifts.
+We consider a simple neural network that predicts reward probability for each (cue card, guide card) pair.
+Two populations, $z_{1}$ and $z_{2}$, encode similarity based on color and shape, respectively.
+A third population, $y=\phi(w_{1}z_{1}+w_{2}z_{2})$, integrates this information to predict the reward probability.
+Here $\phi(x)=\text{sigmoid}(x-\text{bias})$ represents the neural transfer function.
+The weights $w_{1}$ and $w_{2}$ combine the features to determine the prediction, and the guide card with the highest predicted reward probability is selected.
+The agent dynamically updates the model based on received rewards to adapt to changing contexts.
+Traditionally, weights $w_{1}$ and $w_{2}$ are considered synaptic weights, and learning is attributed to synaptic plasticity.
+Alternatively, $w_{1}$ and $w_{2}$ can be reinterpreted as the activity of two additional populations acting as gain modulators In our model, we refer to them as "virtual weights," as they replace traditional physical synaptic weights, effectively virtualizing the learning process.
+These modulatory populations dynamically reconfigure the responses of the network, enabling fast, flexible adaptations without the need for synaptic changes.
+One of the key advantages of this virtualization is that it allows for learning rules that are not constrained by the strict locality requirements inherent to synaptic plasticity.
+To support this point, in Fig.
+1 we compare synaptic learning using the gradient-based, delta rule $\Delta\mathbf{w}=\eta e\cdot\mathbf{z}$ with a non-local learning rule that force constraints $w_{1}(t)+w_{2}(t)=const$.
+This results in the update rule $\Delta w_{j}=\eta(-1)^{j}e\cdot(z_{2}-z_{1})$.
+Here $e(t)=(R(t)-y(t))$ represents the reward prediction error for the last choice and $\eta$ is the learning rate, controlling the speed of adaptation.
+As the latter learning rule features non-local terms, it can easily be implemented by a network that dynamically adjusts the activities of two modulatory populations of virtual weights $w_{1},w_{2}$ without relying on synaptic updates (see Fig.
+1 C).
+In Fig.
+1 we compare the performance of an agent that learns with synaptic plasticity with an agent that performs dynamic adaptation.
+As we can see in panel 1 F, decreasing the timescales of the learning process, the network with dynamic adaptation achieves maximum reward after seeing only one in-context example, while the performance of the network with synaptic plasticity plateaus at near chance level.
+This is supported by the psychometric curve shown in Fig.
+1 G.
+Interestingly, the comparison between the two curves in Fig.
+1 G closely mirrors findings from a recent study [30].
+In that study, some human participants displayed the ability to adjust their behavior after just one example, while others, with impaired cognitive abilities, had to relearn the previous classification rule through trial and error from scratch.
+In this work, we propose biologically plausible framework for understanding rapid behavioral adaptations in dynamic environments, bridging the gap between biological and artificial intelligence.
+Figure 1: Adapting behavior in a context-dependent classification task.
+A.
+Overview of the task structure: At each iteration, the agent is presented with a cue card and two guide cards, one matching the cue by color and the other by shape.
+The agent must select the correct guide card to receive a positive reward.
+The classification rule switches after 20 trials.
+B.
+Schematic representation of the proposed network architecture: synaptic plasticity of physical weights is replaced by dynamic adaptation of virtual weights C.
+that modulate the network response.
+D.
+Example temporal trajectory for the plastic network.
+Two in-context errors are needed to learn the correct rule.
+E.
+Example temporal trajectory for the network with dynamic adaptation: one in-context errors are needed to learn the correct rule.
+F.
+Average reward after 1 in-context example as a function of the learning rate $\eta$.
+Synaptic plasticity with a local delta rule does not go beyond the chance level.
+Dynamic adaptation with a non-local learning rule can achieve a perfect score.
+The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
+The shaded region represents the standard deviation confidence interval.
+G.
+Psychometric curve comparing average reward as a function of in-context examples.
+At $-1$ we report the average reward before the rule switch.
+The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
+Framework Formalization
+Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$.
+The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f$ from the environment in return.
+This can be an error, reward, or the target behavior.
+We further assume that a low-level network extracts features $\mathbf{z}(t)$ relevant to the task family.
+The agent then adapts its response to the specific environment by learning readout parameters $\mathbf{w}$, which determine how to recombine these features to produce the correct response, $y(t)=Y(\mathbf{z}(t),\mathbf{w})$.
+Here $Y$ is a generic function that depends on the parameters $\mathbf{w}$.
+These parameters are then adjusted at each time step in function of an internally computed error signal $e$.
+The overall learning process can then be generically rewritten as the following dynamical system:
+$\begin{cases}\tau_{e}\,\dot{\mathbf{e}}&=E(\mathbf{e},\mathbf{y},\mathbf{f}),\\ \tau_{y}\,\mathbf{y}&=Y(\mathbf{y},\mathbf{w},\mathbf{z}),\\ \tau_{w}\,\dot{\mathbf{w}}&=W(\mathbf{w},\mathbf{z},\mathbf{e}).\end{cases}$ (1)
+The specific instantiation of the functions $W$ and $E$ depends on the learning algorithm employed and can encompass both supervised and reinforcement learning settings.
+Refer to the Methods section for their precise definitions.
+In the classic view, the parameters $\mathbf{w}$ are interpreted as synaptic weights of a neuronal network, and the functions $E$ and $W$ describe synaptic plasticity mechanisms.
+However, this imposes stringent locality requirements, as all information must be present at the synapse.
+Rather than interpreting $\mathbf{w}$ as synaptic weights, we view them as the dynamic activity of an auxiliary network that modulates the response $Y$ in real time based on feedback and past interactions.
+We refer to these dynamically adapting state variables as "virtual" weights.
+We refer to $\mathbf{w}$ as “virtual weights” because: - Weights: they replace the physical weights, and they have the same role as the physical weights of parametrizing the map from the features $\mathbf{z}(t)$ to the output(s).
+- Virtual: they are technically not weights, since the are encoded in the activity of the readout neurons of the network W, and their activity is provided as an input current to the Y network.
+Using this strategy we replace the computation done by the synapses (including learning) with operations performed by the neural activity.
+Dynamic adaptation offers key advantages over synaptic plasticity:
+  • Fast Timescales: It naturally operates on faster timescales, making it well-suited for rapid behavioral changes
+  • Non-Locality: Dynamic adaptation is not restricted by the stringent locality requirements of synaptic plasticity, allowing for richer forms of interaction and computation at the network level.
+In this interpretation, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights.
+We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning.
+While these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the circuit level.
+We demonstrate that such circuits capable of dynamic learning can simply arise by tuning the static readout weights of random RNNs with multiplicative interactions, requiring only a limited number of training trajectories from the original learning algorithm applied to a subset of the task family.
+The construction of our networks follows two distinct phases: a developmental or evolutionary phase, and a subsequent phase where the network leverages this structure for dynamical learning.
+  • Phase 1: Development of the network structure.
+In this phase, using an Algorithmic Distillation protocol[31], we collect learning trajectories of the form $\{(\mathbf{z}_{\alpha}(t),\mathbf{w}_{\alpha}(t),\mathbf{e}_{\alpha}(t),\mathbf{f}_{\alpha}(t))_{t=0}^{T}\|\alpha\in\mathcal{A}_{ID}\}$ by running the learning dynamics in (1) on a subset of tasks $\mathcal{A}_{ID}$ (In Distribution).
+We then train the static readout weights $\Theta$ of two random RNNs $Y_{\Theta_{y}}$ and $W_{\Theta_{w}}$ to replicate, respectively, the functions $Y$ and $W$ on the training data.
+  • Phase 2: The whole architecture is tested (dynamical learning).
+The architecture is tested on a new, previously unseen task $\alpha\in\mathcal{A}_{OOD}=\mathcal{A}\setminus\mathcal{A}_{ID}$ (Out-of-Distribution).
+During this phase, all synaptic weights remain fixed.
+Instead of relying on synaptic plasticity, the virtual weights evolve dynamically according to: $\mathbf{w}(t+dt)=\mathbf{w}(t)+W_{\Theta_{w}}(t)$.
+The updated output of the network is given by $y(t)=Y_{\Theta_{y}}(\{\mathbf{w}(t),\mathbf{z}(t)\})$.
+Importantly, we remark that $\mathbf{w}(t)$ no longer represents physical synaptic weights but rather contextual signals encoded in the activity of neurons, acting as input currents to the $Y_{\Theta_{y}}$ network.
+The role of gain-modulation in adaptive behavior
+Recent research has highlighted a growing interest in understanding modulatory mechanisms and their role in neural computation.
+It has been observed that gain modulation allows neurons to adjust their response to inputs based on context [24].
+This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility.
+Moreover, it has been implicated in various functions, including feedback control [32, 33]) and attention, where it allows for the selective amplification of relevant information.
+A recent study [34] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex.
+Furthermore, gain modulation is closely tied to meta-learning and multitask learning [35, 36], enabling neural networks to efficiently generalize across diverse tasks by leveraging contextual cues.
+Within biological recurrent networks, dendritic segregation provides a structural basis for gain modulation.
+By compartmentalizing synaptic inputs, neurons can effectively implement non-linear interactions between different input streams.
+This is reminiscent of In-Context Learning (ICL) in machine learning, where multiplicative interactions are crucial for enabling models to leverage previously learned information in new contexts.
+Although we focus on gain modulation mediated by apical dendrites, other mechanisms are also known.
+For instance, neuro-glia interactions at the synapse have been proposed as a substrate for working memory, with astrocytes playing a crucial role in modulating neural activity and maintaining memory through glia–synapse interactions [37].
+This gain modulation mechanism could complement other adaptive processes, such as subtle, correlated changes in connectivity [33].
+Our work explores the role of gain modulation in recurrent networks, employing a simplified model where apical currents act multiplicatively on basal currents within a transfer function.
+A key parameter, $\gamma$, controls the strength of these non-linear interactions (see Methods for further details).
+By modulating the gain of neuronal responses, these networks exhibit enhanced adaptability and in-context learning capabilities.
+Results
 Figure 2: Dynamical adaptation of temporal trajectories: A.
 Overview of the network architecture employed.
-A recurrent network composed of $N=20$ units (described by the vector $\bm{z}(t)$, depicted in pink, which follows Eq.
+A recurrent network composed of $N=20$ units (described by the vector $\mathbf{z}(t)$, depicted in pink, which follows Eq.
 (4)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing.
@@
 Dynamical adaptation for temporal trajectories
-We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$.
-Following the reservoir computing paradigm, we employ an RNN (Fig.
-1A, pink), whose dynamics follows Equation (6), to extract temporal features.
-During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x=y^{targ}(t)$ in Eq.
-(6), while $R$ is Gaussian matrix with zero mean and variance, $\sigma_{R}^{2}$) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig.
-1B).
-This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN.
-The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$.
+We consider the task of autonomously predicting a temporal trajectory $\{f(t)=y^{targ}(t)\}_{t}$.
+Following the reservoir computing paradigm (see Methods section), we employ a random RNN (Fig.
+2A, pink), whose dynamics follows Equation (4), to extract temporal features $\mathbf{z}(t)$.
+During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x(t)=y^{targ}(t)$ in Eq.
+(4)) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig.
+2B).
+This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section.
+The readout weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$.
 This results in the following dynamics:
-$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},\bm{w})\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\,e\end{cases}$ (9)
-Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.1A green and orange respectively).
-Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section.
+$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},\mathbf{w}\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},e\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\,e\end{cases}$ (2)
+[…]
 These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$.
 More specifically, the target sequences are $5$ sinusoidal functions with different frequencies (see vertical red lines in Fig.
-1C.)
-In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq.
+2C.)
+In the closed-loop phase, the features $\mathbf{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq.
 (4)).
@@
 (see Fig.
-1B).
+2B).
 In Fig.2B we report an example of successful dynamical adaptation of our model.
@@
 Each curve represents the average regret per round for one fixed trained model, computed by averaging over 100 independent simulations.
-3.1 Multi Armed Bandits
+Extension to Reinforcement learning and the role of Gain modulation
 Dynamical adaptation is now investigated within the framework of reinforcement learning.
@@
 During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
-In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$.
-Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig.
-2A, green).
+In this simplified bandit scenario, where state information is absent, the virtual weights $\mathbf{w}$ directly parameterize the policy probabilities: $\mathbf{\pi}=\mathrm{softmax}{(\mathbf{w})}$.
+Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}^{\gamma}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig.
+3A, green).
 This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
-To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action.
-We train this network to approximate the policy gradient update rule.
-As training data, we use policy gradient estimates over one single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment.
-We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) models in out-of-distribution (OOD) environments.
+The network $W_{\Theta^{w}}^{\gamma}$$(\mathbf{\pi},a,r)$ has a gain-modulated architecture, as described in the Methods section, and is a function of the reward $r$ and the action $a$.
+Here $\gamma$ indicates the dependence on the gain-modulation strength.
 We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment.
@@
 Similar to A, C, E, respectively, but without gain modulation.
-3.2 Reinforcement learning in a reaching task.
-We examine a reaching task, known as the dark room task (a simple instance of the water maze task [31]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$.
+Reinforcement learning in a reaching task.
+We examine a reaching task, known as the dark room task (a simple instance of the water maze task [39]),
+set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$.
 Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position.
@@
 Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility.
-The position of the agent is encoded separately using 25 input units $\bm{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
+The position of the agent is encoded separately using 25 input units $\mathbf{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
 To accomplish this task we consider a network architecture composed of two networks.
-One GM-network, $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig.
-3A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig.
-3A, shown in orange) resulting from changes in virtual weights.
-Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section.
+One GM-network, $W_{\Theta^{w}}^{\gamma}$($(\{\mathbf{z},r,a,\mathbf{\pi}\})$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig.
+4A, green network), while another GM-network $Y_{\Theta^{y}}^{\Gamma}$($(\{\mathbf{z},\mathbf{w}\})$ estimates policy reconfiguration (Fig.
+4A, shown in orange) resulting from changes in virtual weights.
+Here, the parameters $\gamma,\Gamma$ define the strength of the gain modulation, as explained in the Methods section.
 Their dynamics can be described by the following equation Eq.(10).
 Indeed, we compare the case with (Fig.
-3A) and without (Fig.
-3B) gain modulation.
+4A) and without (Fig.
+4B) gain modulation.
 We refer to Supporting Information for further details on the training procedure.
 Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig.
-3C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
-3E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
+4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
+4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
 After multiple trials, the agent successfully achieved precise targeting of the circles.
@@
 To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig.
-3C, left panel) and new test positions (OOD, Fig.
-3C, right panel).
+4C, left panel) and new test positions (OOD, Fig.
+4C, right panel).
 For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training.
@@
 We present the reward plotted against the number of trials for both training (Fig.
-3E, right panel, pink line) and test (Fig.
-3C, right panel, pink line) food locations.
+4E, right panel, pink line) and test (Fig.
+4C, right panel, pink line) food locations.
 We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig.
-3B, D, F).
+4B, D, F).
 Notably, while performances for ID food locations are acceptable (Fig.
-3E, left panel, pink lines), those for OOD cases are extremely poor (Fig.
-3E, right panel, pink lines).
+4E, left panel, pink lines), those for OOD cases are extremely poor (Fig.
+4E, right panel, pink lines).
 This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
-We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section C in Appendix), requiring evaluating temporal credit assignment.
-4 Discussion
-It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 32], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8].
-Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers.
-This explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11].
-However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals.
-In this this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
-When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity.
-Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward.
-We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates.
-Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response.
-We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation.
-Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
-In conclusion, our approach provides an explicit framework to induce ICL in biologically plausible networks, possibly opening the route to a formal understanding of ICL in biological agents.
-Limitations of this study
-In this study, we focused on dynamical adaptation of virtual readout weights, in other words, only the last layer is considered.
-We are neglecting the dynamics on recurrent weights and weights of the hidden units, which, in general, strongly affect performance and stability.
-Generalizability to untested real-world conditions is uncertain, and we do not investigate the scalability to larger architectures.
-We discuss that the mechanism we discuss could be present in biological agents, allowing in-context learning.
-However, the theoretical basis for attention mechanisms executing gradient descent updates lacks comprehensive empirical validation in biological experiments.
-Additionally, the study does not fully address how dynamic gradient structures could be learned or developed in biological networks.
-5 Acknowledgments
-This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’) to MM.
-LF is supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.
+Temporal credit assignment and Future Reward
+We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section Reinforcement Learning and the Discount Factor in Appendix), requiring evaluating temporal credit assignment.
+Indeed, in reinforcement Learning (RL) an agent learns to maximize cumulative rewards.
+The discount factor, usually denoted by $\gamma_{d}$ (where $0\leq\gamma_{d}\leq 1$), is crucial in RL as it determines the importance of future rewards.
+A higher $\gamma_{d}$ values future rewards more significantly, encouraging the agent to consider the long-term consequences of its actions.
+Conversely, a lower $\gamma_{d}$ makes the agent prioritize immediate rewards.
+The cumulative reward $R_{t}$ at time step $t$ is given by:
+$R_{t}=\sum_{s\geq t}r_{s}\gamma^{(s-t)}_{d}$ (3)
+To account for this, the term $\mathbf{e}$ in Eq.(10) should be replaced by its exponential temporal filtering, with a timescale $\tau_{e}=-\frac{1}{\log(\gamma_{d})}$ (see also Eq.(13) in the Appendix).
+In this way actions in the past, can be potentiated when delayed rewards are received.
+If the discount factor is zero, the agent’s policy would change only when it is very close to the moment the reward is received.
+In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points (see Fig.
+A2.D-E).
+As a result, the policy points towards the target position only when nearby the target itself (see Fig.
+A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig.
+A2F, black line).
+On the other hand, recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions, achieving optimal planning and decisions even when far from the target location (and hence, from the reward, see Fig.
+A2.A-B-C).
+Discussion
+The brain’s ability to flexibly and rapidly adapt to novel tasks remains only partially understood.
+Different studies have highlighted that this adaptability relies on mechanisms such as dynamic network reconfiguration and context-dependent modulation, allowing agents to respond flexibly to environmental changes.
+However, a significant gap remains in understanding how these processes map onto specific neural substrates.
+While synaptic plasticity has been extensively studied as a mechanism underlying learning and memory [40], emerging evidence suggests that it alone may not account for the full spectrum of adaptive behavior.
+Abraham et al.
+critically evaluate the synaptic plasticity and memory (SPM) hypothesis, concluding that non-synaptic mechanisms, such as intrinsic plasticity, structural neural changes, epigenetic regulation, glial contributions, network reorganization, and protein turnover, likely contribute significantly to memory storage and retrieval, thus broadening the framework for understanding dynamic adaptation.
+This perspective questions long-standing assumptions about neural circuits, highlighting that we may still lack full understanding of many processes we consider well-established.
+Inspired by the brain’s adaptability, artificial neural network architectures have been challenged to replicate similar capabilities.
+Fast synaptic plasticity, hypothesized to orchestrate flexible cognitive functions, inspired several artificial architectures, such as fast-weight networks [41, 42] and hyper-networks that predict main network weights [43].
+However, these approaches often rely on biologically unrealistic mechanisms, such as requiring neurons to directly act on the synapses of a second network or implementing "inner loops" in network dynamics [44, 43].
+More biologically plausible alternatives, like coupled neural-synaptic dynamics [45], come closer to natural systems but still depend on complex credit assignment mechanisms.
+Transformers, known for their remarkable in-context learning (ICL) capabilities, excel at tasks requiring adaptability, such as natural language and image processing [14, 15, 16].
+Mechanisms such as attention, associative memory [17], and induction-based copying by attention heads [18] have been proposed to explain ICL in transformers, with theoretical work showing that attention mechanisms can implement gradient updates [19].
+However, the emergence of these capabilities remains poorly understood, particularly when comparing them to the in-context learning observed in biological systems.
+We propose a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
+We offer an alternative formulation in which dynamic adaptation is modulated by dendritic input segregation and amplification.
+Virtual (fast) weights are encoded by the activity of a second network that modulates the transfer function of neurons in the base network.
+As a result, virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other neurons, which were pre-trained to perform gradient updates.
+Another pre-trained network receives those weights signals as an input, along with the current environment features, and computes an adapted in-context response.
+This mechanism overcomes the strict locality constraints of synaptic plasticity, which require all elements involved in evaluating weight updates to be present at the synaptic site.
+Consequently, biologically plausible synaptic updates cannot explicitly rely on information from other weights, a requirement often seen in backpropagation methods.
+In contrast, virtual weights are encoded by neuronal activity and can propagate across the neural circuit, bypassing these limitations.
+This allows virtual weight updates to depend on other virtual weights, enabling a more natural and precise evaluation of credit assignment.
+Our work generalizes and provides a biologically plausible implementation for deep gated RNNs performing ICL [46], maintaining a similar number of trainable parameters.
+By dividing the architecture into two components and introducing an additional objective that forces one network to approximate the gradient of the other, we achieve robust dynamic adaptation in tasks requiring temporal learning.
+When tasked to learn temporal trajectories, through error feedback and virtual weight updates, our network achieves successful dynamical adaptation, without synaptic plasticity.
+Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, which are updated in function of the reward.
+We find that networks with gain modulation exhibit improved performance and robustness compared to those without gain modulation.
+Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing their capability to face novel scenarios.
+While we do not fully address how the static weights of the networks involved can be learned using biologically plausible learning rules, we demonstrate that these architectures naturally emerge by tuning the readout weights of random networks.
+These weights could be shaped through evolutionary processes, adapted over slow timescales via reinforcement learning, or optimized using local plasticity rules based on feedback alignment.
+We argue that this represents a step forward in biological plausibility compared to existing deep-learning models like transformers, gated RNNs, and hyper-networks, which rely on more complex and biologically implausible credit assignment mechanisms such as backpropagation through time to train synaptic weights.
+In conclusion, our approach provides an explicit framework to induce in-context learning (ICL) in biologically plausible networks, offering a constructive pathway toward understanding ICL in biological systems and bridging the gap between neuroscience and machine learning.
+Methods
+Gain modulated reservoir computing (GM-RC)
+Reservoir Computing
+Reservoir Computing (RC) represents a paradigm in machine learning that provides a model for biologically plausible computation and learning, drawing inspiration from the information processing mechanisms observed in biological neural networks.
+In this approach, a random RNN is employed to extract features from a time-dependent signal $\mathbf{x}(t)$.
+The dynamics of the RNN describe the evolution of $N$ hidden units $\mathbf{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
+$\tau_{z}\dot{\mathbf{z}}=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}$ (4)
+The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([47]).
+Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively.
+Subsequently, the features extracted by the network can be utilized to predict a target signal $\mathbf{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\mathbf{y}^{targ}(t)-\Theta\mathbf{z}(t)|^{2}$ is minimized.
+The reservoir is then implementing a map
+$\mathbf{y}(t)=\mathrm{RNN}_{\Theta}(\{\mathbf{x}(s)\}_{s\leq t})$ (5)
+where $\Theta$ represents trainable parameters.
+In the rest of the article, we will write $\mathbf{y}=\mathrm{RNN}_{\Theta}(\{\mathbf{x}\})$, dropping the temporal dependencies.
+111 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [48].
+When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit.
+In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\mathbf{x}\right)$.
+Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights.
+This architecture is also referred to in the literature as the Extreme Learning Machine [49].
+Gain modulated network architecture
+Analysing Equations (8) and (10) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters.
+Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC).
+This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [50, 51].
+Here, we consider an additional input source $\mathbf{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$.
+Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope.
+Consequently, the resulting RNN equation is formulated as:
+$\tau_{z}\dot{\mathbf{z}}$ $=\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(J\mathbf{z}+\beta\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}+R\mathbf{x})\right)-\mathbf{z}$ (6)
+Where $\mathbf{b}^{\text{ap}}$ is a constant bias vector.
+The hyperparameters $\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs.
+Specifically, when $\gamma=0$, we obtain a network in which $\mathbf{x}^{\text{ap}}$ does not affect the gain modulation.
+Similarly, as before, the expression $\mathbf{y}=\mathrm{GMRNN}^{\gamma}_{\Theta}(\{\mathbf{x}|\mathbf{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir.
+As before, removing time dependencies, we will have an instantaneous function $\mathbf{y}=\mathrm{GMNN}^{\gamma}_{\Theta}(\mathbf{x}|\mathbf{x}^{\text{ap}})=\Theta\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(R\mathbf{x})\right)$ of the inputs $\mathbf{x}^{\text{ap}}$ and $\mathbf{x}$.
+We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network.
+The choice for the name $\mathbf{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [50].
+Dynamical supervised learning for temporal trajectory
+We consider the task of learning a target temporal trajectory.
+Each environment $\alpha$ is defined by a target trajectory $y_{\alpha}^{targ}(t)$, such that the input and the feedback $x_{\alpha}(t)=f_{\alpha}(t)=y_{\alpha}^{targ}(t)$ are both given by the target sequence.
+The network response is defined as the current estimation of $y^{targ}(t)$ via a linear readout $y(t)=\mathbf{w}\cdot\mathbf{z}(t)$, where $\mathbf{z}$ are the activity of a random RNN that operates as a reservoir computer extracting nonlinear features of the input sequence.
+The RNN dynamics is given by equation (4), with input $x$ given by the target sequence.
+The readout weights in turn dynamically adapt by following the delta rule, minimizing the reconstruction error.
+The learning process can then be described by the following dynamical system:
+$\begin{cases}e&=(y^{targ}-y)\\ y&=\mathbf{z}\cdot\mathbf{w}\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{z}\,e\end{cases}$ (7)
+where we removed the dependence on time $t$ for simplicity.
+The operations required are nonlinear, and usually are viewed as implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights.
+However, here we interpret $\mathbf{w}$ as dynamic variables, proposing, as a possible implementation, that such non-linear functions are computed by two neural networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
+$\begin{cases}e&=(y^{targ}-y)\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{x}|\mathbf{w}\})\simeq Y(\mathbf{x},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{x}|e\})\simeq W(\mathbf{x},e)=\mathbf{x}\,e\end{cases}$ (8)
+In particular, we used two GMRNNs that receive $\mathbf{x},\mathbf{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm).
+Notice that now (8) represents an RNN dynamically performing the learning process.
+Dynamical reinforcement learning
+In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behavior across diverse environments.
+We assume the same state transition dynamics across all the environments $\dot{\mathbf{x}}=X(\mathbf{x},a)$, while the reward distribution $r\sim Prob_{\alpha}(r|\mathbf{x},a)$ depends on the current environment $\alpha\in\mathcal{A}$.
+Here $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones.
+We start by defining a policy network, denoted as $\mathbf{\pi}=\mathrm{softmax}(\mathbf{y})$ which implements a policy mapping the agent state encoded by the feature vector $\mathbf{z}$, to a probability distribution over actions.
+We assume an agent that linearly recombines these features $\mathbf{y}=Y(\mathbf{z},\mathbf{w})=\mathbf{w}\cdot\mathbf{z}$.
+This assumption is made without loss of generality, since we could always employ a reservoir computer as an intermediate feature extraction layer, as described in the previous subsection.
+The agent learns by evaluating the policy gradient with respect to the weights $\mathbf{w}$.
+This leads to a dynamics similar to equation (8), where the definition of $e(t)$ is changed as follows:
+$\begin{cases}\mathbf{e}&=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}\right)\\ \mathbf{y}&=\mathbf{w}\cdot\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{e}\odot\mathbf{z}\end{cases}$ (9)
+Here $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [52, 53]) at time $t$.
+It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
+This formulation holds for a null discount factor [52, 53]), we refer to the Appendix section Reinforcement Learning and the Discount Factor for the description of the general case.
+As before, we model that the policy is modulated by an auxiliary population of virtual weights $\mathbf{w}$, determined by the activity of an additional GMRNN $W_{\Theta_{y}}$.
+This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behavior in real-time.
+Then, the above dynamics can be approximated by two networks, one for estimating the gradients, and one for the scalar product.
+$\begin{cases}\tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{z}|r,a,\mathbf{\pi}\})\simeq W(\mathbf{z},r,a,\mathbf{\pi})=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}(\mathbf{y})\right)\odot\mathbf{z}\\ \mathbf{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{z}|\mathbf{w}\})\simeq Y(\mathbf{z},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\end{cases}$ (10)
+Acknowledgments
+This research has received financial support from the Italian National Recovery and Resilience Plan (PNRR), M4C2, funded by the European Union - NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’).
+LF aknowledges support by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing and Sapienza University of Rome (AR12419078A2D6F9).
+Work partially funded by the European Union – NextGenerationEU and the Ministry of University and Research (MUR), Italian National Recovery and Resilience Plan (PNRR), M4C2I1.3, project ‘MNESYS’ (PE00000006).
+We sincerely thank Maurizio Mattia for helpful discussions on the topic.
 Source code availability
 The source code is available under CC-BY license in the https://github.com/cristianocapone/ABSS public repository.
-Appendix A Relationship with previous work on product-based architectures implementing in context-learning
-The relationship between attention-based transformer architectures and in-context learning was first noted in [11] where it was shown through constructive proof, confirmed by experiments, that linear attention layers can implement a gradient descent update for linear regression in its forward pass.
-Building upon this insight, [14] showed that a linear gated RNN can implement the same mechanism, showing that a linear two-layer gated RNN can replicate a linear transformer.
-The work proposes an implementation that uses $O(d^{2})$ hidden units and has $O(d^{4})$ trainable weights, that can be reduced to $O(d^{3})$ by side gating.
-Our gain modulated architecture significant analogies with the gated RNNs analyzed in [14].
-Consider a linear  [^3] For simplicity, here we assume $\alpha,\beta=0,\ \gamma=1$ and consider a linear activation function $\phi=Id$.
-gain-modulated network architectures implementing the functions ${\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})$, ${\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)$, dynamically performing a ICL linear regression task that requires to map temporal dependent features $z(t)\in\mathbb{R}^{N_{in}}$ to predict an output $\bm{y}^{targ}\in\mathbb{R}^{N_{out}}$:
-$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})=\Theta^{y}\left(R^{ap}_{y}\bm{w}\right)\odot(R_{y}\bm{z}(t))\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},\bm{y}^{targ}-\bm{y})=\Theta^{w}\left(R^{ap}_{w}\left(\hat{y}-y\right)\right)\odot(R_{w}\bm{z}(t))\end{cases}$ (10)
-This architecture can be seen as a particular instance of the RNN with side gating (see Equation (15) in [14]).
-Interestingly, a similar mechanism is also present in the recently proposed Mamba layer [33].
-Appendix B Approximating gradients with gain modulated architectures
-In this appendix, we show that an architecture with gain modulation is more suited to approximate gradient terms involved in dynamical learning.
-Scalar product
-We first consider the task of approximating a scalar product between virtual weights $\bm{w}\in\mathbb{R}^{N_{in}}$ and features $\bm{x}\in\mathbb{R}^{N_{in}}$.
-To achieve this, we train the readout weights $\Theta$ of a gain modulated network $\mathrm{GMNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{w}\})$ with $N_{h}$ hidden features to approximate the function $dot(\bm{x},\bm{w})=\sum_{i=1}^{N_{in}}x_{i}w_{i}$.
-The training dataset is composed of $1000$ pairs $(\bm{x},\bm{w})$ uniformly sampled in the hypercube $[0,1]^{2N_{in}}$.
-The test set consists of the same number of pairs sampled in the hypercube $[-1,1]^{2N_{in}}$.
-We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
-For each of the two settings, we select the best hyperparameters by fixing the dimensionality of the inputs ($N_{in}=5$) and hidden units ($N_{h}=101$) and varying the standard deviation of the projection matrices $R_{ij},R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}),\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
-The best hyperparameters found were used in the models to test the approximation performance of the dot product varying the number of input features ($N_{in}\in[1,10]\cap\mathbb{N}$) and hidden units ($N_{h}\in[1,200]\cap\mathbb{N}$).
-Results are shown in Fig.
-A1 A, B, C.
-We see that while both architectures require a quadratic number of hidden units to achieve low error on the test set, the gain-modulated network after a threshold number of units is hidden units is reached, is able to perfectly approximate the scalar product function consistently achieving $\sim 10^{-7}$ RMSE error.
-At the same time, the error in the network without gain modulation remains several orders of magnitude higher.
-Scalar-vector product
-We then tackle the task of approximating a vector-scalar product between features $\bm{x}\in\mathbb{R}^{N_{in}}$ and features and error $\bm{e}\in\mathbb{R}$.
-To achieve this, we train the readout weights $\Theta$ of a gain modulated network $\mathrm{GMNN}_{\Theta}(\{\bm{x}\}|\{\gamma e\})$ with $N_{h}$ hidden features to approximate the function $prod(\bm{x},e)=e\bm{x}\in\mathbb{R}^{N_{in}}$.
-The training dataset is composed of $1000$ pairs $(\bm{x},e)$ uniformly sampled in the hypercube $[0,1]^{N_{in}}\times[0,1]$.
-The test set consists of the same number of pairs sampled in the hypercube $[-1,1]^{N_{in}}\times[-1,1]$.
-We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
-For each of the two settings, we select the best hyperparameters by fixing the dimensionality of the inputs ($N_{in}=5$) and hidden units ($N_{h}=21$) and varying the standard deviation of the projection matrices $R_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}/N_{in}),R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2},\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
-The best hyperparameters found were used in the models to test the approximation performance of the dot product varying the number of input features ($N_{in}\in[1,10]\cap\mathbb{N}$) and hidden units ($N_{h}\in[1,40]\cap\mathbb{N}$).
-Results are shown in Fig.
-A1 D, E, F.
-We see that the network without gain modulation requires a quadratic number of hidden units to achieve low error on the test set.
-As in the scalar product case, the gain-modulated network can nearly perfectly approximate the target function after a threshold number of hidden units is reached.
-Significantly, in this case, the threshold scales linearly with $\mathbb{R}^{N_{in}}$.
-Figure A1: A Test errors for the scalar product approximation task, varying the number of features $N_{in}$, hidden units $N_{h}$.
-We compare architectures with gain modulation ($\gamma=1$, bottom) with architectures without gain modulation ($\gamma=0$, bottom).
-B Test errors for the dot product approximation task, varying the number of features $N_{in}$, and fixing the number of hidden units $N_{h}=101$.
-Solid lines indicate the median over 100 trained models, while the filled region indicates the $20/80th$-percentile interval.
-C Same as B but fixing the number of features $N_{in}=5$, and varying the number of hidden units $N_{h}=101$.
-D, E, F Same as panels A, B, C but for the scalar-vector product approximation task.
-On the number of hidden units and trainable weights needed in our model
-Consider the architecture in Eq. (10) with $N_{out}=1$:
-$\begin{cases}y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})=\Theta^{y}\left(R^{ap}_{y}\bm{w}\right)\odot(R_{y}\bm{z}(t))\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)=\Theta^{w}\left(R^{ap}_{w}\left(\hat{y}-y\right)\right)\odot(R_{w}\bm{z}(t))\end{cases}$ (11)
-Where $R^{ap}_{y},R_{y},R^{ap}_{w},R_{w}$ are fixed random matrices respectively of dimension $N_{y}\times N_{in},N_{y}\times N_{in},N_{w}\times 1,N_{w}\times N_{in}$, and $\Theta^{y},\Theta^{w}$ are trainable readout matrices of sizes $1\times N_{y},N_{in}\times N_{w}$.
-In this simplified linear setting, the features extracted by the network are linear combinations of the product of the input entries that need to be linearly combined by the readout weights to the target function, which in both cases is composed of a specific linear combination of these products.
-It is then straightforward to observe that, to approximate the dot product, in ${\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{z},\bm{w})$ we need $N_{w}\sim O(N_{in}^{2})$ hidden units and $O(N_{in}^{2})$ readout parameters.
-as confirmed in the experiment.
-This number can be further reduced to $O(N_{in})$ both for the hidden units and trainable weights  [^4] This can be done assuming that $R^{ap},R_{w}$ are diagonal matrices (and thus reducing the problem to estimating $N_{in}$ $1\times 1$ products).
-As for the network ${\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{z},y^{targ}-y)$ approximating vector-scalar products, similar considerations support the experimental observation that $N_{y}\sim O(N_{in})$ hidden units and $O(N_{in}^{2})$ features are needed in this case.
-For a multivariate $\mathbb{R}^{N_{in}}\to{N_{out}}$ linear regression task, we can consider $N_{out}$ independent modules of the type described before.
-In this case, the number of trainable parameters is $O(N^{2}_{in}\cdot N_{out})$ with $O(N_{in}\cdot N_{out})$ hidden units ($O(N_{in}^{2}\cdot N_{out})$ in the more biologically plausible case).
-Appendix C Reinforcement Learning and the Discount Factor
-Reinforcement Learning (RL) is a machine learning paradigm where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards.
-The discount factor, usually denoted by $\gamma$ (where $0\leq\gamma\leq 1$), is crucial in RL as it determines the importance of future rewards.
-The cumulative reward $R_{t}$ at time step $t$ is given by:
-$R_{t}=r_{t}+\gamma r_{t+1}+\gamma^{2}r_{t+2}+\gamma^{3}r_{t+3}+\ldots$
-In our experiment, the policy gradient update rule in the presence of the discount factor could be approximated as (see [23]):
-$\begin{cases}\bm{e}&=\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=r\bm{\hat{e}}\end{cases}$ (12)
-where $\bm{\hat{e}}$ is an exponential temporal filtering of $\bm{e}$, with a timescale that is proportional to $-\frac{1}{\log(\gamma)}$.
-The importance of the discount factor lies in its ability to balance immediate and future rewards.
-A higher $\gamma$ values future rewards more significantly, encouraging the agent to consider the long-term consequences of its actions.
-Conversely, a lower $\gamma$ makes the agent prioritize immediate rewards.
-This balance is essential for the stability of the learning process.
-An appropriately chosen $\gamma$ ensures stable learning and convergence; if $\gamma$ is too high, the agent might overvalue distant future rewards, leading to instability, while a very low $\gamma$ can result in short-sighted behavior.
-If the discount factor is zero, the agent’s policy would change only when it is very close to the moment the reward is received.
-In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points.
-To demonstrate that our approach performs the temporal computation required to consider future rewards, we compare the performances of our network (see Fig.
-A2.A-B) to a network without recurrent connections, which therefore cannot perform temporal computations.
-Recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions.
-Without these connections, a network operates purely on the current input without any contextual information from prior steps, thus lacking the ability to perform temporal computations (see Fig.
-A2.D-E).
-Performances are higher in the first case.
-This can be visualized by looking at the policy at the end of the dynamic reinforcement learning for a specific target location.
-When the recurrent weights are set to zero, the policy points towards the target position only when nearby the target itself (see Fig.
-A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig.
-A2F, black line).
-On the other hand, in the presence of recurrent weights, the proper policy (pointing towards the target, see Fig.
-A2C) is known even when far from the target, allowing optimal long-term planning (see Fig.
-A2C, black line).
-Figure A2: Discount Factor and eligibility traces: A.
-Sample trajectories of the agent at the end of the learning procedure, for various target locations (color-coded), different from the ones used to train the gradient and the scalar product networks (red crosses).
-B.
-Reward as a function of number of trials (or games), blue: policy gradient, pink: dynamical learning for new positions (OOD).
-Line: median, shading: 20-th/80-th percentile range.
-C.
-Arrows: policy at the end of the training as a function of the position.
-Line: sample trajectory of the agent at the end of learning.
-Small circle: agent position, double circle: target position.
-D-E-F same as in A-B-C, but the recurrent connections of the gradient network are set to zero.
-Appendix D Additional details on the "dynamical learning of a temporal trajectory" experiment
-In the "dynamical learning of a temporal trajectory" experiment we test our architecture and non-synaptic learning approach on temporal tasks.
-The primary goal is to analyze the network’s ability to generalize beyond the target frequencies used to pre-train our networks.
-Table 1: Simulation Parameters
-Parameter Symbol Reservoir Gradient Net Scalar Net
-Network Size $N$ 20 500 500
-Input Dimension $I$ 1 100 + 10 100+100
-Apical Input Dimension $I^{ap}$ 0 1 100
-Output Dimension $O$ 1 100 1
-Time Step $dt$ 0.005
-Reservoir Time Constant $\tau_{m_{f}}$ 10 $dt$ 1 $dt$ 1 $dt$
-Input weights var $\sigma_{input}$ 0.06 0.06 0.06
-Apical Input weights var $\sigma_{input}^{ap}$ 0.0 0.1 0.1
-Recurrent weights $\sigma_{rec}$ 0.99 / $\sqrt{N}$ 0.5 / $\sqrt{N}$ 0. / $\sqrt{N}$
-Gain-modulation factor $\gamma_{net}$ 0. 1. 1.
-Three networks are defined in this experiment: one uses a reservoir to compute temporal features and encode the target temporal trajectory, one predicts the required weight updates, and one predicts the scalar product, as discussed in the main text.
-The parameters used in the simulation are summarized in Table 1.
-Inputs are projected to the network through Gaussian weights with zero mean and variance $\sigma_{in}^{2}$ to distribute the input information across multiple units in the reservoir.
-Data collection and pre-training
-The readout of the first reservoir network is defined as $R_{j}$, and the readout at time $t$ is given by:
-$y(t)=\sum_{j}R_{j}z_{j}(t),$
-where $z_{j}(t)$ are the reservoir states.
-Training and evaluation involve presenting the networks with five target trajectories $y^{targ}(t)=0.8sin(\omega_{targ}t)$, with five angular velocities, $\omega_{targ}$, ranging from $0.04$ to $0.08$.
-The reservoir readout parameters are trained using online gradient descent to minimize the error between the current output and the target output.
-The loss function is defined as:
-$L=\frac{1}{2}\left(y(t)-y^{\text{target}}(t)\right)^{2}.$
-The gradient used for online training is:
-$\Delta R_{j}=-\eta\frac{\partial L}{\partial R_{j}}=-\eta\left(y(t)-y^{\text{target}}(t)\right)z_{j}(t),$
-where $\eta$ is the learning rate.
-During data collection, trajectories of network states, errors, and weight updates are recorded.
-The gradient and scalar-product networks are trained on this data to predict the required gradient given a specific temporal error, and to estimate the current trajectory given the current virtual weights.
-The networks are pre-trained by estimating only their readout weights, in accordance with a reservoir computing prescription. They are trained using a pseudo-inverse approach.
-Dynamical supervised online learning and evaluation
-Once the architecture is pre-trained, it can be tested to learn new trajectory, unobserved during the pretraining.
-readout weights of the reservoir are no longer changed online using gradient, but rather virtual weights are changed following the prescription of the gradient network.
-Virtual weights are then used as an input to the scalar-product network to predict the current prediction $y(t)$.
-Performances are evaluated by measuring the MSE between the target trajectory ant the predicted one $y(t)$, for different values of the trajectory angular velocity, equally distribute between 0.01.01.01.01 and $0.1$.
-Appendix E Additional details on the bandit experiment
-Family of tasks
-We consider a family of tasks $\mathcal{A}$, such that every element $\alpha\in\mathcal{A}$ represents a Bernoulli $K$-armed bandit problem with rewards $\{R_{i}^{\alpha}\}_{i=1}^{K}$.
-Each task $\alpha$ is specified by a set $P_{\alpha}\subset\{1\cdot,K\}$ of positive arms, such that
-$R_{i}^{\alpha}=\begin{cases}Ber({p})&i\in{P_{\alpha}}\\ Ber({1-p})&i\in{P_{\alpha}}\end{cases}$ (13)
-In the main text experiment, we used $K=10$ and $P_{ID}:=\{i\in[K]|i\equiv 0(\text{mod}2)\}$ and $P_{OOD}:=\{i\in[K]|i\equiv 1(\text{mod}2)\}$.
-Network details and hyperparameter search
-We compare an architecture with gain modulation ($\gamma=1$) with an architecture without gain modulation $\gamma=0$.
-For each of the two settings, we select the best hyperparameters by fixing the number of hidden units ($N_{h}=100$) and varying the standard deviation of the projection matrices $R_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}/20)\ \ R^{ap}_{ij}\sim\mathcal{N}(0,\sigma_{R}^{2}),\ \log_{10}(\sigma_{R})\in\{-2,-1.8,\cdots,0\}$, the standard deviation of the bias $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2}),\ \sigma_{b}\in\{0,0.1,\cdots,1\}$ and the nonlinearity type $\phi\in\{\mathrm{tanh},\mathrm{softplus}\}$.
-For each point in the grid, we train 20 models using The best hyperparameters found were used in the subsequent experiments In Fig.
-2 B, D the reported regrets are on models trained with $N_{h}=100$ hidden features
-Regret per round
-Given an agent that plays in an environment $\alpha$ receiving rewards $\{r_{t}\}_{t\in\mathbb{N}_{+}}$, the regret per round $\rho(T)$ achieved by the agent at round $T$ is defined as:
-$\rho(T)=\mu^{\alpha}_{\star}-\frac{1}{T}\sum_{t=1}^{T}r_{t}$ (14)
-Where $\mu^{\alpha}_{*}:=\max\{\mathbb{E}[R_{k}^{\alpha}]|k\in[K]\}$ is the maximum expected reward that can be obtained in each round playing the optimal policy that selects one of the best arms with probability $1$.
-Appendix F Additional details on the darkroom experiment
-This experiment investigates the capability of out dynamical reinforcement learning approach, to improve an agent’s ability to navigate a 2D environment.
-The agent is trained to locate a randomly placed, not observable food source, with its policy evolving over multiple episodes through.
-The environment is a 2D grid where an agent starts at the center, aiming to reach a randomly positioned food item.
-The food positions vary every 600 episodes, introducing.
-The agent’s movements are restricted within the grid, with actions limited to moving left, right, up, or down.
-To represent the agent’s position, a place cell encoder converts the agent’s $(x,y)$ coordinates into a higher-dimensional feature vector using Gaussian functions.
-This encoding helps in effectively capturing spatial information.
-The agent’s policy is linear, with action probabilities derived from a softmax function applied to the encoded state.
-For this reason, only two networks are used for the task, the gradient and the scalar-product network see Table 2 for details on parameters.
-Data collection and pre-training
-The policy is updated using the policy gradient method, adjusting weights based on received rewards to improve decision-making over time.
-Learning data are collected for 8 different positions of the food, equally distributed on a circle.
-Two types of networks are used: one to model the agent’s state dynamics and another to handle gradient updates necessary for learning.
-These networks influence the agent’s internal state and learning process, enabling policy refinement.
-The two networks readout are trained with a linear regression on their readout weight, to reproduce the proper weight updates an scalar products.
-Dynamical reinforcement learning and evaluation
-The agent is tested again on the 8 positions used for pre-training, and on 8 new positions.
-Similarly to previous experiments, policy gradient is no longer used and replaced by virtual weights updated estimated by the gradient network, and used by the scalar-product network to predict agent policy.
-Performance is measured by the total reward accumulated over episodes. Visualizing the agent’s trajectories reveals its navigation efficiency and decision-making process.
-Table 2: Simulation Parameters
-Parameter Symbol Gradient Net Scalar Net
-Network Size $N$ 500 1000
-Input Dimension $I$ 25+4+10 4 $\times$ 25+4 $\times$ 25
-Apical Input Dimension $I^{ap}$ 1 4 $\times$ 25
-Output Dimension $O$ 4 $\times$ 25 4
-Time Step $dt$ 0.005
-Reservoir Time Constant $\tau_{m_{f}}$ 1 $dt$ 1 $dt$
-Input weights var $\sigma_{input}$ 0.01 0.01
-Apical Input weights var $\sigma_{input}$ 0.1 0.1
-Recurrent weights $\sigma_{rec}$ 0.5 / $\sqrt{N}$ 0. / $\sqrt{N}$
-Gain-modulation factor $\gamma_{net}$ 1. 1.

### v5

@@
 We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning.
-Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
+We further validate our approach in a realistic robotic setting, demonstrating dynamic adaptation in a MuJoCo ant navigation task without synaptic updates, showcasing a neuromorphic control paradigm that operates through real-time network reconfiguration.
+Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence, and provides a measurable approach to distinguish between dynamic adaptation and synaptic plasticity, aiding experimental validation.
 Introduction
@@
 This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
-Main / Framework description
-Dynamical adaptation replaces synaptic plasticity: A popular example
-Consider a task inspired by the Wisconsin card-sorting test.
-Each card in a deck is defined by two features: color and shape.
-In each iteration, the agent receives a cue card and two guide cards—one matching the cue by color and the other by shape.
-The agent must select the correct guide card to receive a positive reward $R(t)=1$.
-Initially, the correct choice is based on color matching.
-However, after 10 trials, the classification rule changes, rewarding shape matching instead.
-A successful agent must quickly adapt its decision-making to these contextual shifts.
-We consider a simple neural network that predicts reward probability for each (cue card, guide card) pair.
-Two populations, $z_{1}$ and $z_{2}$, encode similarity based on color and shape, respectively.
-A third population, $y=\phi(w_{1}z_{1}+w_{2}z_{2})$, integrates this information to predict the reward probability.
-Here $\phi(x)=\text{sigmoid}(x-\text{bias})$ represents the neural transfer function.
-The weights $w_{1}$ and $w_{2}$ combine the features to determine the prediction, and the guide card with the highest predicted reward probability is selected.
-The agent dynamically updates the model based on received rewards to adapt to changing contexts.
-Traditionally, weights $w_{1}$ and $w_{2}$ are considered synaptic weights, and learning is attributed to synaptic plasticity.
-Alternatively, $w_{1}$ and $w_{2}$ can be reinterpreted as the activity of two additional populations acting as gain modulators In our model, we refer to them as "virtual weights," as they replace traditional physical synaptic weights, effectively virtualizing the learning process.
-These modulatory populations dynamically reconfigure the responses of the network, enabling fast, flexible adaptations without the need for synaptic changes.
-One of the key advantages of this virtualization is that it allows for learning rules that are not constrained by the strict locality requirements inherent to synaptic plasticity.
-To support this point, in Fig.
-1 we compare synaptic learning using the gradient-based, delta rule $\Delta\mathbf{w}=\eta e\cdot\mathbf{z}$ with a non-local learning rule that force constraints $w_{1}(t)+w_{2}(t)=const$.
-This results in the update rule $\Delta w_{j}=\eta(-1)^{j}e\cdot(z_{2}-z_{1})$.
-Here $e(t)=(R(t)-y(t))$ represents the reward prediction error for the last choice and $\eta$ is the learning rate, controlling the speed of adaptation.
-As the latter learning rule features non-local terms, it can easily be implemented by a network that dynamically adjusts the activities of two modulatory populations of virtual weights $w_{1},w_{2}$ without relying on synaptic updates (see Fig.
-1 C).
-In Fig.
-1 we compare the performance of an agent that learns with synaptic plasticity with an agent that performs dynamic adaptation.
-As we can see in panel 1 F, decreasing the timescales of the learning process, the network with dynamic adaptation achieves maximum reward after seeing only one in-context example, while the performance of the network with synaptic plasticity plateaus at near chance level.
-This is supported by the psychometric curve shown in Fig.
-1 G.
-Interestingly, the comparison between the two curves in Fig.
-1 G closely mirrors findings from a recent study [30].
-In that study, some human participants displayed the ability to adjust their behavior after just one example, while others, with impaired cognitive abilities, had to relearn the previous classification rule through trial and error from scratch.
-In this work, we propose biologically plausible framework for understanding rapid behavioral adaptations in dynamic environments, bridging the gap between biological and artificial intelligence.
-Figure 1: Adapting behavior in a context-dependent classification task.
-A.
-Overview of the task structure: At each iteration, the agent is presented with a cue card and two guide cards, one matching the cue by color and the other by shape.
-The agent must select the correct guide card to receive a positive reward.
-The classification rule switches after 20 trials.
-B.
-Schematic representation of the proposed network architecture: synaptic plasticity of physical weights is replaced by dynamic adaptation of virtual weights C.
-that modulate the network response.
-D.
-Example temporal trajectory for the plastic network.
-Two in-context errors are needed to learn the correct rule.
-E.
-Example temporal trajectory for the network with dynamic adaptation: one in-context errors are needed to learn the correct rule.
-F.
-Average reward after 1 in-context example as a function of the learning rate $\eta$.
-Synaptic plasticity with a local delta rule does not go beyond the chance level.
-Dynamic adaptation with a non-local learning rule can achieve a perfect score.
-The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
-The shaded region represents the standard deviation confidence interval.
-G.
-Psychometric curve comparing average reward as a function of in-context examples.
-At $-1$ we report the average reward before the rule switch.
-The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
+Main
 Framework Formalization
 Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$.
-The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f$ from the environment in return.
+The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f(t)$ from the environment in return.
 This can be an error, reward, or the target behavior.
@@
   • Non-Locality: Dynamic adaptation is not restricted by the stringent locality requirements of synaptic plasticity, allowing for richer forms of interaction and computation at the network level.
-In this interpretation, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights.
+In the Supplementary Materials, we show that a non-local weight update rule, implemented through dynamic adaptation, enables rule switching with a single error feedback in a context-dependent classification task, whereas local synaptic plasticity drops to chance level after one error—independently of the timescales of the learning process (Fig.
+A2).
+In the dynamic adaptation setting, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights.
 We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning.
@@
 This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility.
-Moreover, it has been implicated in various functions, including feedback control [32, 33]) and attention, where it allows for the selective amplification of relevant information.
+Moreover, it has been implicated in various functions, including feedback control [31, 32] and attention, where it allows for the selective amplification of relevant information.
 A recent study [33] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex.
@@
 Following the reservoir computing paradigm (see Methods section), we employ a random RNN (Fig.
-2A, pink), whose dynamics follows Equation (4), to extract temporal features $\mathbf{z}(t)$.
+1A, pink), whose dynamics follows Equation (4), to extract temporal features $\mathbf{z}(t)$.
 During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x(t)=y^{targ}(t)$ in Eq.
 (4)) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig.
-2B).
+1B).
 This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section.
@@
 More specifically, the target sequences are $5$ sinusoidal functions with different frequencies (see vertical red lines in Fig.
-2C.)
+1C.)
 In the closed-loop phase, the features $\mathbf{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq.
@@
 (see Fig.
-2B).
+1B).
 In Fig.1B we report an example of successful dynamical adaptation of our model.
@@
 Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}^{\gamma}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig.
-3A, green).
+2A, green).
 This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
@@
 After several trials, the agent achieves precise targeting of the circles.
-Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel).
 C.
@@
 Similar to A, C, E, respectively, but without gain modulation.
-Reinforcement learning in a reaching task.
-We examine a reaching task, known as the dark room task (a simple instance of the water maze task [39]),
-set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$.
+Dynamical Reinforcement learning in a reaching task
+We examine a reaching task, known as the dark room task (a simple instance of the water maze task [38]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$.
 Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position.
@@
 One GM-network, $W_{\Theta^{w}}^{\gamma}$($(\{\mathbf{z},r,a,\mathbf{\pi}\})$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig.
-4A, green network), while another GM-network $Y_{\Theta^{y}}^{\Gamma}$($(\{\mathbf{z},\mathbf{w}\})$ estimates policy reconfiguration (Fig.
-4A, shown in orange) resulting from changes in virtual weights.
+3A, green network), while another GM-network $Y_{\Theta^{y}}^{\Gamma}$($(\{\mathbf{z},\mathbf{w}\})$ estimates policy reconfiguration (Fig.
+3A, shown in orange) resulting from changes in virtual weights.
 Here, the parameters $\gamma,\Gamma$ define the strength of the gain modulation, as explained in the Methods section.
@@
 Indeed, we compare the case with (Fig.
-4A) and without (Fig.
-4B) gain modulation.
+3A) and without (Fig.
+3B) gain modulation.
 We refer to Supporting Information for further details on the training procedure.
 Firstly, we assessed the performance of the policy gradient algorithm using a set of 8 food locations (refer to Fig.
-4C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
-4E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
+3C, indicated by red crosses), where the total reward averaged over 200 trials was observed against the number of trials (Fig.
+3E, depicted by blue lines, thin lines for individual positions and thick lines for the average across all positions).
 After multiple trials, the agent successfully achieved precise targeting of the circles.
@@
 To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig.
-4C, left panel) and new test positions (OOD, Fig.
-4C, right panel).
+3C, left panel) and new test positions (OOD, Fig.
+3C, right panel).
 For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training.
@@
 We present the reward plotted against the number of trials for both training (Fig.
-4E, right panel, pink line) and test (Fig.
-4C, right panel, pink line) food locations.
+3E, right panel, pink line) and test (Fig.
+3C, right panel, pink line) food locations.
 We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig.
-4B, D, F).
+3B, D, F).
 Notably, while performances for ID food locations are acceptable (Fig.
-4E, left panel, pink lines), those for OOD cases are extremely poor (Fig.
-4E, right panel, pink lines).
+3E, left panel, pink lines), those for OOD cases are extremely poor (Fig.
+3E, right panel, pink lines).
 This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
@@
 In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points (see Fig.
-A2.D-E).
+A3.D-E).
 As a result, the policy points towards the target position only when nearby the target itself (see Fig.
-A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig.
-A2F, black line).
+A3F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig.
+A3F, black line).
 On the other hand, recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions, achieving optimal planning and decisions even when far from the target location (and hence, from the reward, see Fig.
-A2.A-B-C).
+A3.A-B-C).
+Application to a realistic task
+We test our framework to a more complex and realistic setting by applying it to a MuJoCo ant environment, demonstrating its scalability to continuous control tasks.
+The agent, modeled as a quadruped, operates in a high-dimensional control space with 27 environmental state variables and 8 control torques (2 per leg), requiring coordinated control of multiple degrees of freedom.
+To manage low-level motor execution, we first trained a network to perform four fundamental motor primitives—moving left, right, forward, and backward.
+This pretraining step allows higher-level adaptation without direct joint-level optimization (see Appendix for details).
+Building on this, we use our gain-modulated network as a meta-policy that dynamically selects motor primitives based on the agent’s position.
+The meta-policy network (orange network Fig.
+4A) receives positional inputs and determines the appropriate primitive, while a secondary network (green network Fig.
+4A) modulates its behavior through virtual weights, mirroring the architecture used in the grid-based task.
+This architecture enables flexible and efficient adaptation: the agent successfully generalizes to target locations not included in the training set (see Fig.
+4C), adjusting its movement strategies dynamically without modifying its underlying synaptic structure.
+This capability is particularly significant for neuromorphic computing and robotics, where adaptive behavior is often constrained by the need for on-chip plasticity.
+Our results demonstrate that a fixed neuromorphic chip, programmed with a gain-modulated network, can dynamically adjust to new goals without requiring structural modifications.
+This approach reduces computational overhead and power consumption while enabling efficient, low-power autonomous systems that adapt in real time.
+By eliminating the need for weight updates at the hardware level, this architecture provides a scalable and robust foundation for adaptive physical agents, paving the way for next-generation neuromorphic robotics.
+Figure 4: Dynamical Reinforcement Learning Dark Room A.
+Overview of the network architecture employed.
+One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights.
+B.
+Illustration of the task, and of the mojoco environment: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles).
+After several trials, the agent achieves precise targeting of the circles.
+Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel).
+C.
+Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD).
+Line: median, shading: 20-th/80-th percentile range.
 Discussion
@@
 Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing their capability to face novel scenarios.
-While we do not fully address how the static weights of the networks involved can be learned using biologically plausible learning rules, we demonstrate that these architectures naturally emerge by tuning the readout weights of random networks.
-These weights could be shaped through evolutionary processes, adapted over slow timescales via reinforcement learning, or optimized using local plasticity rules based on feedback alignment.
-We argue that this represents a step forward in biological plausibility compared to existing deep-learning models like transformers, gated RNNs, and hyper-networks, which rely on more complex and biologically implausible credit assignment mechanisms such as backpropagation through time to train synaptic weights.
+Our findings demonstrate that adaptive network architectures can emerge by selecting appropriate readout weights in initially random networks, without requiring explicit backpropagation.
+These weights can be shaped through biologically plausible mechanisms such as evolutionary processes, slow reinforcement learning, or local plasticity rules relying on feedback alignment.
+Unlike standard deep-learning models—such as transformers, gated RNNs, and hyper-networks—that depend on non-local credit assignment mechanisms like backpropagation through time, our approach leverages local, dynamically reconfigurable processes, making it more compatible with biological learning principles.
+In line with this, our model suggests that dynamic reconfiguration provides a biologically plausible mechanism for rapid adaptation, complementing traditional fast synaptic plasticity.
+Different neural circuits in the brain may rely on varying balances of dynamic reconfiguration and synaptic plasticity, shaped by their functional roles and evolutionary pressures.
+In the Supplementary Materials, we provide a proof of concept through a T-maze simulated experiment, where we observe a substantial difference in how contextual information is encoded for synaptic versus adaptive learning.
+This distinction supports the potential for experimental validation.
 In conclusion, our approach provides an explicit framework to induce in-context learning (ICL) in biologically plausible networks, offering a constructive pathway toward understanding ICL in biological systems and bridging the gap between neuroscience and machine learning.